### v1

# Generating variable $\hbar$ and $c$ via Fujii–Wetterich model in curved spacetimes

Abstract
We revisit the Fujii–Wetterich model [Phys.Rev.D 26, 2580 (1982) and Nucl.Phys.B 302, 645 (1988)] which allows the Higgs doublet to couple with a ‘cosmon’ scalar $\chi$ of the background spacetime as $\chi^{2}\,\Phi^{2}$. Upon spontaneous breaking of the $SU(2)$ gauge symmetry, the vacuum expectation value (VEV) of the Higgs doublet is proportional to the ‘cosmon’ field $\chi$. Fujii and Wetterich employed this linkage to make particle masses dependent on $\chi$. We shall present an alternative mechanism: at a given point $x^{*}$, the prevailing Higgs VEV will be used to construct a quantum of action $\hbar_{*}$ and a speed of light $c_{*}$ in association with $\chi(x^{*})$. Specifically, each open set vicinity of a point $x^{*}$ on the manifold is equipped with a replica of the Glashow–Weinberg–Salam action operating with its own effective values of $\hbar_{*}$ and $c_{*}$, whereas particle masses induced via Higgs spontaneous symmetry breaking remain independent of $\chi(x^{*})$. Our mechanism unambiguously generates the dependencies $\hbar_{*}\propto\chi^{-1/2}(x^{*})$ and $c_{*}\propto\chi^{1/2}(x^{*})$, causing these ‘fundamental constants’ to vary along with the dynamical field $\chi$ across the manifold. Moreover, in each open set around $x^{*}$, the field $\chi(x^{*})$ determines the length and time scales for physical processes occurring in this region as follows: $l\propto\chi^{-1}(x^{*})$ and $\tau\propto\chi^{-3/2}(x^{*})$. The resulting anomalous anisotropic relation between the rate of clocks and the length of rods, $\tau^{-1}\propto l^{-3/2}$, leads to a concrete physical prediction, manifesting as a new time dilation effect, and produces a distinct set of previously unknown phenomenology.
For late-time cosmology, the variation of $c$ along the trajectory of light waves from distant supernovae towards the Earth-based observer renders the classic Lemaître redshift formula $1+z=a^{-1}$ inapplicable. Using the dependency $c_{*}\propto\chi^{1/2}(x^{*})$, we derive the new (variable-$c$) Lemaître redshift relation and apply it to analyze the Pantheon Catalog of Type Ia supernovae without invoking the dark energy hypothesis. Detailed expositions of our analysis are provided in companion papers, with the key consequences summarized herein: (1) Accounting for the Pantheon Catalog with a fit exceeding the quality of the $\Lambda$CDM model; (2) Explaining the late-time cosmic acceleration based solely on variable $c$, eliminating the need for dark energy; (3) Revitalizing Blanchard–Douspis–Rowan-Robinson–Sarkar’s CMB power spectrum analysis that bypassed dark energy [A&A 412, 35 (2003) and A&A 449, 925 (2006)]; (4) Resolving the age problem without resorting to dark energy; and (5) Resolving the $H_{0}$ tension through the evolution of galaxies on an expanding cosmic background, rather than requiring a dynamical dark energy component.
I Theoretical consideration
Motivation—In the light of several outstanding issues and tensions in cosmology, there has emerged growing interest in promoting the parameters of physical models to be variable scalar fields in spacetime. One popular example is the dynamical dark energy which could offer relief to the fine tuning and coincidence problems in late-time cosmology. The cosmological constant $\Lambda$ is equipped with a kinetic term that allows it to evolve and ‘relax’ to its (small) value in the current epoch.
Against this backdrop, the accustomed Planck $cGh$ unit system (1899-1900) solidifies the role of the speed light $c$, the (Newton) gravitational constant $G$, and the quantum of action $\hbar$ as “fundamental constants” which encompass the realms of special relativity, gravitation, and quantum mechanics, respectively [1, 2]. The trio $\{c,\,G,\,\hbar\}$ are viewed as fixed cornerstones, convertible to the three “fundamental units”—the Planck mass, the Planck length, and the Planck time, defined as
$m_{P}:=\sqrt{\frac{\hbar c}{G}};\ \ l_{P}:=\sqrt{\frac{\hbar G}{c^{3}}};\ \ \tau_{P}:=\sqrt{\frac{\hbar G}{c^{5}}}\,.$ (1)
These units span the basis for describing all physical phenomena known to date.
This orthodox view has faced challenges, however. The most well-known example is the concept of variable $G$, inspired by Dirac in 1937 [3], and culminating in the generally covariant Brans-Dicke (BD) theory of gravity in 1961 [4]. In this theory, a dynamical scalar degree of freedom $\chi$ is introduced alongside the metric tensor $g_{\mu\nu}$ as
$\int d^{4}x\,\frac{\sqrt{-g}}{16\pi}\left[\chi^{2}\,\mathcal{R}-4\omega\,\nabla_{\mu}\chi\nabla^{\mu}\chi\right]$ (2)
where $\omega$ is a dimensionless (BD) parameter. The BD theory has grown into a family of scalar–tensor theories, a very popular theme of research nowadays. The field $\chi^{2}$, often known a ‘dilaton’, also naturally arises in string theory and other contexts [5].
On the face value, the (Newton) gravitational ‘constant’ becomes a scalar field via its dependency on the field $\chi$, via a formal identification
$\frac{1}{16\pi}\chi^{2}\,\mathcal{R}:=\frac{c^{3}}{16\pi\hbar G}\,\mathcal{R}\ \ \ \Longrightarrow\ \ \ G\propto\chi^{-2}\,.$ (3)
With the aid of Eq. (1), this dependency results in
$m_{P}\propto\chi;\ \ \ l_{P}\propto\chi^{-1};\ \ \ \tau_{P}\propto\chi^{-1}$ (4)
In the vanilla BD theory, the matter sector coupled minimally to the gravitational sector to preserve Einstein’s Equivalence Principle [4]. However, around the same time, Utiyama and DeWitt showed that even if one starts with a minimal coupling action between matter and gravity, radiative corrections would generate non-minimal coupling between them [6]. Since the 1980s, efforts to introduce non-minimal coupling directly at tree level have been pursued independently in the works of Fujii and Wetterich, which we shall call collectively as the Fujii–Wetterich scheme. In [7] Fujii allowed the Higgs doublet $\Phi$ of the Glashow–Weinberg–Salam (GWS) model of particle physics to couple with the Ricci scalar in the form $\mathcal{R}\,\Phi^{2}$. In [8, 9, 10, 11, 12] Wetterich embedded the BD action in a broader framework that projects a dilatation symmetry, with the non-minimal coupling in the form  [^1] Wetterich referred the field $\chi$ a ‘cosmon’ [8]. In place of Eq. (5), he considered a more general form $V(\chi,\Phi)=\Phi^{4}\,w\left(\chi^{2}/\Phi^{2}\right)$ where $w$ is a function of the ratio $x:=\chi^{2}/\Phi^{2}$. If $w(x)$ adopts an affine form $\lambda-\mu^{2}\,x$, then Eq. (5) would ensue.
$V(\chi,\Phi)=\lambda\,\Phi^{4}-\mu^{2}\,\chi^{2}\,\Phi^{2}$ (5)
where $\lambda\in\mathbb{R^{+}}$ and $\mu\in\mathbb{R}$ are dimensionless parameters.
An essential element in the Fujii–Wetterich scheme is that, when the $SU(2)$ gauge symmetry of the Higgs doublet undergoes spontaneous breaking by way of the potential prescribed in Eq. (5), the field $\chi$ influences the vacuum expectation value (VEV) of the Higgs doublet $\Phi$, which in turn affects the mass of the fermions and the gauge vector bosons in the GWS model. Specifically, Fujii and Wetterich demonstrated that the electron mass $m$ scales proportional to $\chi$, thereby maintaining a constant ratio $m/m_{P}$. Moreover, the length scale $l$ and time scale $\tau$ of any given physical process also become dependent on $\chi$ as $l\propto\chi^{-1}$ and $\tau\propto\chi^{-1}$, in exact proportion to $l_{P}$ and $\tau_{P}$ respectively. Most importantly, the field $\chi$ affects not only the Planck time per Eq. (4) but also the rate of the clocks (viz. the revolution rate of a mechanical clock, the oscillation rate of an atomic clock, or the decay rate of an unstable quantum system) which are made of matter governed by the GWS model. Quantitatively, the Fujii–Wetterich scheme predicts that the evolution rate $\tau^{-1}$ of physical processes is proportional to $\chi$.  [^2] We find it illuminating to quote Fujii [7]: “… the time and length in the microscopic unit frame are measured in units of $m^{-1}(t)$, in agreement with the physical situation that the time scale of atomic clocks, for example, is provided by the atomic levels which are determined by the Rydberg constant $(me^{4})^{-1}$ and Wetterich [10]: “The clock provided by the Hubble expansion in the standard description is now replaced by a clock associated to the increasing value of $\chi$.
It is imperative to stress that the field $\chi$ itself serves as a (dynamical) length scale that can be used for constructing $G$, a dimensionful quantity. It is then logically permissible to adopt an alternative mechanism whereby the field $\chi$, a quantity equal to $\sqrt{c^{3}/(\hbar G)}$, could also be used to derive $c$ and $\hbar$, rather than $G$. The key objective of our paper is to establish the dependencies of $\hbar$ and $c$ on $\chi$ via the matter sector. Our approach constitutes a major departure from the Fujii–Wetterich scheme: by requiring the charge and mass of elementary particles to be independent of $\chi$, we unambiguously obtain $c\propto\chi^{1/2}$ and $\hbar\propto\chi^{-1/2}$ while $G$ remains constant. Furthermore, via the time evolution of quantum states $i\hbar\frac{\partial}{\partial t}\left|\psi\right\rangle=\hat{H}\left|\psi\right\rangle$, the dependency $\hbar\propto\chi^{-1/2}$, along with $\hat{H}\propto\chi$, causes the clock rate of physical processes to scale as $\tau^{-1}\propto\chi^{3/2}$, in decisive distinction from the Fujii–Wetterich scheme (which posits $\tau^{-1}\propto\chi$).
The anisotropic time scaling enabled in our approach, characterized by its anomalous $3/2$–exponent, leads to new physics with a distinct set of phenomenology andpredictions. We will discuss these significant implications in this paper and its companions [13, 14, 15, 16].
History of variable $c$ and $\hbar$—Prior to the Brans-Dicke theory, Dicke also briefly considered variability in the speed of light $c$ in 1957 [17] although he did not pursue it further. A relatively little-known fact is that the concept of variable speed of light (VSL) was initially explored by Einstein in 1911, published in three obscure papers (originally in German) [18, 19, 20]  [^3] It seems that Dicke in 1957 [17] was not aware of Einstein’s 1911 VSL attempt [18].. During his quest for a generally covariant theory of gravity (which later became known as the General Relativity), Einstein emphasized that the Michelson–Morley experimental results and Lorentz symmetry are meant to hold only locally. That is to say, while $c$ is invariant with respect to local boosts, its value needs not be universal. Einstein envisioned the possibility that gravitational potential affected the clock rate, hence generating a variable $c$ in spacetime. Such a variation in $c$ could then produce a refraction effect on light ray around the Sun’s disc as a result of Huygens’s principle [18]. The later successes of General Relativity quickly overshadowed Einstein’s 1911 paper on VSL, leading him to abandon the idea, however. In the 1990s, VSL was revived independently by Moffat and by Albrecht and Magueijo to tackle issues in early-time cosmology, such as the horizon paradox [21, 22].
The concept of variable quantum of action $\hbar$ has been much less explored, with a rare exception of [23]. One important empirical guidance is that, although the fine-structure constant $\alpha:=e^{2}/(\hbar c)$ is known to ‘run’ in the renormalization group (RG) flow of quantum field theory, there is no clear experimental evidence supporting $\alpha$ varying in spacetime  [^4] In [24] Webb et al. offered observational evidence of a varying $\alpha$ with respect to the redshift, but this result has been in contention; e.g., see a recent paper [25].. If $\alpha$ is to remain independent of the field $\chi$ while $c$ is allowed to vary, it would require a co-variation of $e$, $\hbar$, and $c$. One such scenario, as briefly mentioned earlier, posits $\hbar\propto\chi^{-1/2}$ and $c\propto\chi^{1/2}$, which would insulate $\alpha$ from depending on $\chi$.
A Higgs–scalon coupling model—Consider the full action
$\mathcal{S}$ $=\int d^{4}x\,\sqrt{-g}\,\mathcal{L}_{\text{mat}}+\int d^{4}x\,\frac{\sqrt{-g}}{16\pi}\,\mathcal{L}_{\text{grav}}$ (6)
in which the matter sector is specified as
$\mathcal{L}_{\text{mat}}$ $=i\,\bar{\psi}\gamma^{\mu}(\nabla_{\mu}-i\sqrt{\alpha}A_{\mu})\psi+f\,\bar{\psi}\psi\,\Phi$
$+\frac{\mu^{2}}{2}\,\chi^{2}\,\Phi^{2}-\frac{\lambda}{4}\,\Phi^{4}+g^{\mu\nu}\partial_{\mu}\Phi\,\partial_{\nu}\Phi-\frac{1}{4}\,F_{\mu\nu}F^{\mu\nu}$ (7)
where all parameters $\alpha$, $f$, $\mu$, and $\lambda$ are dimensionless. The matter sector comprises a massless spinor field $\psi$ coupled with a massless $U(1)$ gauge field $A_{\mu}$, and a charge-neutral Higgs singlet $\Phi$ coupled with a scalar singlet $\chi$ on a 4-dimensional Riemannian spacetime manifold. We shall term the field $\chi$ a ‘scalon’ (a concept to be clarified later), representing a scalar degree of freedom within the gravitational sector.
It is important to remark that the adimensional action of the matter sector does not involve the (dimensionful) quantum of action $\hbar$ and speed of light $c$ at its outset. The length dimensions of $\{\psi,\,A_{\mu},\,\Phi$} are $\{-3/2,-1,-1\}$ respectively. The gamma matrices satisfy $\{\gamma^{\mu}(x),\gamma^{\nu}(x)\}=2\,g^{\mu\nu}(x)$, and the spacetime covariant derivative $\nabla_{\mu}$ acts on the spinor via vierbein and spin connection.
The Lagrangian $\mathcal{L}_{\text{mat}}$ in Eq. (7) is invariant under the local $U(1)$ gauge transformation:
$\psi(x)$ $\rightarrow e^{i\sqrt{\alpha}\sigma(x)}\,\psi(x)$ (8)
$A_{\mu}(x)$ $\rightarrow A_{\mu}(x)-\partial_{\mu}\sigma(x)$ (9)
where $\sigma(x)$ is an arbitrary scalar function. In addition, $\mathcal{L}_{\text{mat}}$ is also invariant under the discrete $\mathbb{Z}_{2}$ symmetry of the Higgs singlet
$\Phi\rightarrow-\Phi\,.$ (10)
Many extended theories of gravitation naturally host a scalar degree of freedom [26]. Regarding $\mathcal{L}_{\text{grav}}$, several possibilities exist, one being a massive Brans-Dicke action
$\mathcal{L}_{\text{grav}}=\chi^{2}\,\mathcal{R}-4\omega\,\nabla^{\mu}\chi\nabla_{\mu}\chi-V(\chi)\,.$ (11)
Since the matter sector in Eq. (7) has a dilatation symmetry (with $\chi$ having the length dimension $-1$), requiring the gravitational sector to also possess this symmetry would force $V(\chi)$ to take the form $\chi^{4}$. When $\omega=0$, $\mathcal{L}_{\text{grav}}$ in Eq. (11) belongs to the $f(\mathcal{R})$ family of gravity [27, 28]. It can also be embedded within a broader family of classically scale invariant gravity (or “agravity”) which includes the Weyl term, $\mathcal{C}^{\mu\nu\lambda\rho}\,\mathcal{C}_{\mu\nu\lambda\rho}$ [30, 29]. Alternatively, it can represent Horndeski gravity, the most general theory in four dimensions constructed out of the metric tensor and a scalar field that leads to second-order equations of motion [31]. Another viable option would be Lyra geometry [32], where the inherent scale function of this geometry can play the role of the scalon field $\chi$. However, the exact details of the gravitational sector are not the focus of our paper. Our work concerns the mattersector, and we require only the existence of a ‘scalon’ singlet $\chi$ that couples with the Higgs field $\Phi$ as described in Eq. (7).
While several choices for $\mathcal{L}_{\text{grav}}$ are available, we shall adopt the pure $\mathcal{R}^{2}$ theory, viz. $\omega=0$ and $V(\chi)\simeq\chi^{4}$, for the gravitational sector in the rest of this paper
$\mathcal{L}_{\text{R2}}=\chi^{2}\,\mathcal{R}-f_{0}\,\chi^{4}\,.$ (12)
At the classical level, this yields $\mathcal{R}=2f_{0}\,\chi^{2}$. Being adimensional, it has dilatation invariance, a crucial feature that proves very convenient for addressing issues in late-time cosmology, as demonstrated in this paper and its companion [14].  [^5] For pure $\mathcal{R}^{2}$ gravity, the Newtonian limit has been established [33, 34]. The theory is also known to be Ostrogradsky stable and free of ghosts [34].
A new mechanism to generate variable $\hbar$ and $c$—As a component of the gravitational sector, the scalon $\chi$ serves as a background field for the matter sector. At a given point $x^{*}$ on the manifold, its value is $\chi_{*}:=\chi(x^{*})$. In the open set vicinity of the point $x^{*}$, the terms $\frac{\mu^{2}}{2}\,\chi_{*}^{2}\,\Phi^{2}-\frac{\lambda}{4}\,\Phi^{4}$ in Eq. (7), evaluated for the prevailing $\chi_{*}$, induce a spontaneous breaking of the discrete $\mathbb{Z}_{2}$ symmetry of the Higgs singlet, given by Eq. (10)  [^6] Note that the $U(1)$ gauge symmetry described by Eqs. (8)–(9) remains unbroken in this process..This process, traditionally known as the Higgs mechanism [35], results in a non-zero Higgs vacuum expectation value (VEV)
$\left\langle\Phi(x^{*})\right\rangle=\frac{\mu}{\sqrt{\lambda}}\chi_{*}$ (13)
which is proportional to $\chi_{*}$. In a local reference frame tangent to the manifold at the point $x^{*}$, the action for the matter sector (excluding the Higgs excitation above the vacuum) becomes
$\int d^{4}x\,\biggl{[}i\bar{\psi}\gamma^{\mu}\partial_{\mu}\psi+\sqrt{\alpha}\,\bar{\psi}\gamma^{\mu}A_{\mu}\psi+\frac{f\mu}{\sqrt{\lambda}}\,\chi_{*}\,\bar{\psi}\psi-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}\biggr{]}\,.$ (14)
Varying the spinor field and the gauge vector field, we obtain two equations of motion
$\left(i\gamma^{\mu}\partial_{\mu}+\sqrt{\alpha}\,\gamma^{\mu}A_{\mu}+\frac{f\mu}{\sqrt{\lambda}}\,\chi_{*}\right)\psi$ $=0$ (15)
and
$\partial_{\nu}F^{\nu\mu}=j^{\mu}:=\sqrt{\alpha}\,\bar{\psi}\gamma^{\mu}\psi\,.$ (16)
The first equation resembles the Dirac equation for the ‘electron’ $\psi$, coupled with a $U(1)$ gauge vector $A_{\mu}$
$\left(i\gamma^{\mu}\partial_{\mu}+\frac{e}{\sqrt{\hbar c}}\,\gamma^{\mu}A_{\mu}+m\,\frac{c}{\hbar}\right)\psi=0\,.$ (17)
The second equation resembles the Maxwell equation for $A_{\mu}$ sourced by a $U(1)$–charged current $j^{\mu}$
$\partial_{\nu}F^{\nu\mu}=j^{\mu}:=\frac{e}{\sqrt{\hbar c}}\,\bar{\psi}\gamma^{\mu}\psi\,.$ (18)
In Eq. (17), the parameters $e$ and $m$ represent the $U(1)$ gauge charge and the mass of the electron, respectively. As they are intrinsic properties of the electron, we then require them to be parameters rather than fields  [^7] We must emphasize that $e$ and $m$ can ‘run’ in the RG flow when radiative corrections related to the matter action described by Eq. (14) are included.. That is to say, we condition that they are independent of $\chi_{*}$. By comparing Eqs. (15)–(16) against Eqs. (17)–(18), we identify
$m:=\frac{f\mu}{\sqrt{\lambda}};\ \ \ e:=\sqrt{\alpha}\,.$ (19)
These identifications then require that
$\hbar\,c:=1;\ \ \ \frac{c}{\hbar}:=\chi_{*}$ (20)
which unambiguously yield the following relations
$\hbar_{*}:=\chi_{*}^{-1/2};\ \ \ c_{*}:=\chi_{*}^{1/2}\,.$ (21)
Here, the subscript * is used to indicate the dependence of $\hbar$ and $c$ as functions of $\chi$. It should be noted that the (Newton) gravitational constant in independent of $\chi_{*}$ because $G:=\frac{c_{*}^{3}}{\hbar_{*}\chi_{*}^{2}}=1$.
Conversely, in the tangent frame covering an open set around the point $x^{*}$, the action in Eq. (14) can be recast as an action of Quantum Electrodynamics (QED) with a Lagrangian expressed in terms of $\hbar_{*}$, $c_{*}$, $m$, and $e$:
$\mathcal{L}_{\text{QED}}=i\bar{\psi}\gamma^{\mu}\partial_{\mu}\psi+\frac{e}{\sqrt{\hbar_{*}c_{*}}}\bar{\psi}\gamma^{\mu}A_{\mu}\psi+m\frac{c_{*}}{\hbar_{*}}\bar{\psi}\psi-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}\,.$ (22)
Therefore, each open set vicinity of a given point $x^{*}$ on the manifold is equipped with a replica of QED action (22), operating with an effective quantum of action $h_{*}$ and an effective speed of light $c_{*}$, both of which are determined by the prevailing value $\chi_{*}$ of the background scalon field. As a component of the gravitational sector, $\chi_{*}$ can vary across the manifold, leading to corresponding variations in $\hbar_{*}$ and $c_{*}$ throughout the manifold.
In each such open set, the effective speed of light $c_{*}$ governs the propagation of the (massless  [^8] Note that only the discrete $\mathbb{Z}_{2}$ symmetry (10) of the Higgs singlet is (spontaneously) broken. The $U(1)$ gauge symmetry in Eqs. (8) and (9) remains unbroken, leaving the gauge vector boson $A_{\mu}$ massless.) gauge vector field $A_{\mu}$, whereas the effective quantum of action $\hbar_{*}$ regulates the quantization of the fields $\psi$, $A_{\mu}$, and $\Phi$. We should note that our mechanism to generate variable $\hbar$ and $c$ is readily applicable to the generalization of $\mathcal{L}_{\text{mat}}$ in Eq. (7) to the full GWS model of particle physics, where the gauge group is enlarged to $SU(3)\times SU(2)\times U(1)$. This extension is relatively straightforward and will be presented elsewhere [13].
The role of $\hbar_{*}$ is instrumental: from Eq. (15) (and using $\hbar_{*}c_{*}=1$ per Eq. (21)), it can be shown that the time evolution of the electron wavefunction is given by
$i\hbar_{*}\frac{\partial}{\partial t}\psi(t)$ $=\hat{H}(t)\,\psi(t)$ (23)
$\hat{H}(t)$ $:=-i\alpha^{k}\frac{\partial}{\partial x^{k}}+\sqrt{\alpha}\left(\alpha^{k}A_{k}-\beta A_{0}\right)+\frac{f\mu}{\sqrt{\lambda}}\,\chi_{*}\,\beta\,.$ (24)
Here, the matrices are $\alpha^{k}:=\gamma^{0}\gamma^{k}$ and $\beta:=\gamma^{0}$. The time evolution of $\psi(t)$ within the open set vicinity of $x^{*}$ thus depends on the behavior of $\hbar_{*}$ with regard to $\chi_{*}$. This crucial connection between $\hbar_{*}$ and the evolution rate of quantum states leads to a concrete prediction, which we shall elaborate in subsequent sections.
A prediction: Anisotropic scaling in the clock rate—The QED action given in Eq. (14) possesses a dilatation symmetry, meaning it remains invariant under the transformation
$dx^{\mu}\rightarrow\chi_{*}\,dx^{\mu};\ \psi\rightarrow\chi_{*}^{-3/2}\psi;\ A_{\mu}\rightarrow\chi_{*}^{-1}A_{\mu}\,.$ (25)
The inverse of the prevailing value $\chi_{*}^{-1}$ of the scalon field $\chi$ at the point $x^{*}$ thus sets the length scale for physical processes within the open set surrounding $x^{*}$, i.e.
$l\propto\chi_{*}^{-1}\,.$ (26)
This is why we have referred to the field $\chi$ as a ‘scalon’.
However, due to the dependence of the speed of light on $\chi_{*}$ as derived in Eq. (21), the time scale for physical processes in the open set surrounding $x^{*}$ exhibits an anomalous behavior
$\tau:=\frac{\chi_{*}^{-1}}{c_{*}}\propto\chi_{*}^{-3/2}\,.$ (27)
This behavior can also be understood through the time evolution operator (23)–(24). Since $dx^{k}\propto\chi_{*}^{-1}$ and $A_{\mu}\propto\chi_{*}$, the Hamiltonian in (24) scales as $\chi_{*}$. The rate of evolution is thus
$\tau\simeq\frac{\hbar_{*}}{\hat{H}}\propto\chi_{*}^{-3/2}$ (28)
which is compatible with the anomalous time scaling found in Eq. (27).
To verify the behavior of the length and time scales with respect to $\chi_{*}$, we shall consider the Hydrogen atom as an example. The Bohr radius is
$a_{B}$ $=\frac{\hbar_{*}}{\alpha\,mc_{*}}=\frac{\sqrt{\lambda}}{\alpha f\mu}\,\chi_{*}^{-1}\propto\chi_{*}^{-1}$ (29)
which is consistent with Eq. (26). The (relativistic) energy level of electron in a quantum state $\left|n,\,j\right\rangle$ of a hydrogen atom is a well-established result [36]  [^9] In the taylor expansion $E_{n}^{j}=mc_{*}^{2}-\frac{\alpha^{2}}{2n^{2}}mc_{*}^{2}+\mathcal{O}\left(\alpha^{4}\right)$, the leading term is the rest-mass energy and the sub-leading term recovers the energy level known in the non-relativistic limit for the Hydrogen atom.
$E_{n}^{j}$ $=\mathcal{N}_{n}^{j}\,mc_{*}^{2}=\mathcal{N}_{n}^{j}\,\frac{f\mu}{\sqrt{\lambda}}\,\chi_{*}$ (30)
in which $\mathcal{N}_{n}^{j}\!:=\!{\scriptstyle{\scriptstyle\left(1+\alpha^{2}\left(n-j-\frac{1}{2}+\sqrt{\left(j+\frac{1}{2}\right)^{2}-\alpha^{2}}\right)^{-2}\right)^{-1/2}}}$ with $n=1,2,3,...$ and $j=\frac{1}{2},\frac{3}{2},\frac{5}{2},...$. Note that energy is proportional to $\chi_{*}$. The groundstate $\left|n=1,\,j=1/2\right\rangle$ and the excited state $\left|n=2,\,j=3/2\right\rangle$ have energy
$E_{n=1}^{j=1/2}$ $=\sqrt{1-\alpha^{2}}\,\frac{f\mu}{\sqrt{\lambda}}\,\chi_{*}$ (31)
$E_{n=2}^{j=3/2}$ $=\sqrt{1-\text{\scriptsize$\frac{1}{4}$}\alpha^{2}}\,\frac{f\mu}{\sqrt{\lambda}}\,\chi_{*}\,.$ (32)
A transition from the (initial) excited state $\left|i\right\rangle=\left|n=2,\,j=3/2\right\rangle$ to the (final) groundstate $\left|f\right\rangle=\left|n=1,\,j=1/2\right\rangle$, induced by electric dipole, is allowed as it satisfies the selection rule $\Delta j=\pm 1$. The energy of the photon emitted is $E_{n=2}^{j=3/2}-E_{n=1}^{j=1/2}$, and the frequency of the emitted photon is
$\nu$ $=\frac{E_{n=2}^{j=3/2}-E_{n=1}^{j=1/2}}{\hbar_{*}}$ (33)
$=\frac{f\mu}{\sqrt{\lambda}}\Bigl{(}\sqrt{1-\text{\scriptsize$\frac{1}{4}$}\alpha^{2}}-\sqrt{1-\alpha^{2}}\Bigr{)}\,\chi_{*}^{3/2}\,.$ (34)
Therefore, the propagation of the photon has the time scale that behaves as
$\tau:=\frac{1}{\nu}\propto\chi_{*}^{-3/2}$ (35)
which is in perfect alignment with Eq. (27). This time scaling result is generic since the time scale of the electron in a quantum state $\left|n,\,j\right\rangle$ can be deduced via Ehrenfest’s theorem, $\left\langle v_{n}^{j}\right\rangle=\frac{1}{m}\left\langle p_{n}^{j}\right\rangle=\frac{1}{mc_{*}}\left(\left(E_{n}^{j}\right)^{2}-m^{2}c_{*}^{4}\right)^{1/2}$. For example, considering the groundstate of the hydrogen atom, the typical time scale for the electron to complete its ‘orbit’ is
$\tau=\frac{2\pi\,a_{B}}{\left\langle v_{n=1}^{j=1/2}\right\rangle}=\frac{2\pi\sqrt{\lambda}}{\alpha^{2}f\mu}\,\chi_{*}^{-3/2}$ (36)
which is in full conformity with Eqs. (27) and (35).
The time scaling (27) implies that the evolution rate of a clock—regardless of whether it is a mechanical clock, an electronic clock, or an atomic clock—varies in spacetime, as a function of $\chi$, in an anisotropic fashion. In principle, this effect can be measured experimentally: Prepare two identical clocks at a location $A$. Keep one clock at location $A$ and send the second clock to a location $B$. Suppose that the background scalon field has different values $\chi_{A}$ and $\chi_{B}$ at the two locations $A$ and $B$, respectively. At their respective locations, the clocks would run at different rates per
$\tau_{A}\propto\chi_{A}^{-3/2};\ \ \ \tau_{B}\propto\chi_{B}^{-3/2}\,.$ (37)
When the clock from location $B$ is brought back to location $A$, it will show a different elapsed time compared to the clock that resided at location $A$ during the whole experiment.
This predicted time dilation effect, while similar to the time dilation effect in GR which is associated with the $g_{00}$ component of the spacetime metric, arises from the dependence of the clock rate on the scalon field $\chi$, viz. Eq. (37). This new phenomenon is distinct and shall be referred to as a “time dilation effect of the Third kind”, to be investigated elsewhere [16].  [^10] In addition to the time dilation effect in GR, there is a well-known effect in Special Relativity where two twice-intersecting time-like paths can have different total amounts of proper time in between. Therefore, we refer our predicted time dilation effect as “the Third kind”.
Decay rate due to spontaneous emission—The prediction that we have just made can be validated with a different setup. Reconsider the Hydrogen atom. Induced by electric dipole perturbation in the vacuum, the electron in the excited state $\left|n=2,\,j=3/2\right\rangle$ can spontaneously transition to the groundstate $\left|n=1,\,j=1/2\right\rangle$ (allowable by the selection rule $\Delta j=\pm 1$). According to quantum mechanics [37], the decay rate (i.e. Einstein’s coefficient) is known to be
$A=\frac{4}{3}\frac{\omega_{if}^{3}e^{2}\left\langle\vec{r}_{if}\right\rangle^{2}}{\hbar_{*}c_{*}^{3}}$ (38)
where we have made the subscripts * in $\hbar$ and $c$ explicit. The angular frequency (with $h_{*}:=2\pi\hbar_{*}$) is
$\omega_{if}=\frac{1}{h_{*}}\left(E_{n=2}^{j=3/2}-E_{n=1}^{j=1/2}\right)$ (39)
and the matrix element of the electric dipole between the initial state $\left|i\right\rangle:=\left|n=2,\,j=3/2\right\rangle$ and the final state $\left|f\right\rangle:=\left|n=1,\,j=1/2\right\rangle$ is $\left\langle\vec{r}_{if}\right\rangle:=\left\langle f|\vec{r}|i\right\rangle$. Using $e^{2}=\alpha\,\hbar_{*}c_{*}$, we get
$A=\frac{4\alpha}{3}\frac{m^{3}c_{*}^{4}}{h_{*}^{3}}\left\langle\vec{r}_{if}\right\rangle^{2}\,.$ (40)
Given that $\left\langle\vec{r}_{if}\right\rangle\propto\chi_{*}^{-1}$, $c_{*}\propto\chi_{*}^{1/2}$, $h_{*}\propto\chi_{*}^{-1/2}$, we find
$A\propto\frac{\chi_{*}^{2}}{\chi_{*}^{-3/2}}\,\chi_{*}^{-2}\propto\chi_{*}^{3/2}\,.$ (41)
Thus, the lifetime of this unstable quantum system behaves as
$\tau:=\frac{1}{A}\propto\chi_{*}^{-3/2}$ (42)
which perfectly conforms with Eq. (27). Consequently, the clock involved in our experiment proposed in the preceding section can be any generic timekeeping device. It can be a mechanical clock, an electronic clock, an atomic clock  [^11] The Hafele–Keating experiment [38, 39] was carried out using atomic clocks., or even a radioactive quantum system.
Comparing with the Fujii–Wetterich scheme—We must stress that the dependence of the clock rate on $\chi_{*}$ has been documented in the works of Fujii and Wetterich [7, 8, 9, 10, 11, 12], although it was not a focal point in their analysis; see Footnote 2. However, their findings diverge from ours, particularly in how the clock rate scales with $\chi_{*}$. Specifically, they predicted a relation
$\tau_{\text{FW}}^{-1}\propto\chi_{*}$ (43)
where “FW” refers to Fujii–Wetterich. This contrasts with our prediction, Eq. (27)
$\tau^{-1}\propto\chi_{*}^{3/2}\,.$
The distinction arises due to a different set of conditions used in Fujii and Wetterich’s treatments [7, 8, 9, 10, 11, 12]. These authors judiciously kept $\hbar$ and $c$ independent of $\chi_{*}$, setting $\hbar_{\text{FW}}=1$ and $c_{\text{FW}}=1$. Instead, they allowed the electron mass $m$ to be a field, while the electron charge $e$ remained a constant. With regard to Eqs. (15)–(16) and Eqs. (17)–(18), their conditions then produce the following identities
$m_{\text{FW}}:=\frac{f\mu}{\sqrt{\lambda}}\,\chi_{*};\ \ \ e:=\sqrt{\alpha}\,.$ (44)
Consequently, the gravitational constant varies as $G_{\text{FW}}:=\frac{c_{\text{FW}}^{3}}{\hbar_{\text{FW}}\chi_{*}^{2}}=\chi_{*}^{-2}$. Instead of Eq. (23), the time evolution of the electron wavefunction in their scheme is governed by
$i\,\hbar_{FW}\,\frac{\partial}{\partial t}\psi(t)$ $=\hat{H}(t)\,\psi(t)$ (45)
with the Hamiltonian $\hat{H}(t)$ given in Eq. (24). Here, $\hbar_{FW}$ is constant (equal to $1$). Therefore, the time scale for evolution in their scheme scales as
$\tau_{\text{FW}}\simeq\frac{\hbar_{\text{FW}}}{\hat{H}}\propto\chi_{*}^{-1}\,.$ (46)
This result can be verified with the concrete examples we used previously. In the Fujii–Wetterich scheme, the frequency of emitted photon from the excited state $\left|n=2,\,j=3/2\right\rangle$ to the groundstate $\left|n=1,\,j=1/2\right\rangle$ is
$\nu_{\text{FW}}$ $=\frac{E_{n=2}^{j=3/2}-E_{n=1}^{j=1/2}}{\hbar_{\text{FW}}}$ (47)
$=\frac{f\mu}{\sqrt{\lambda}}\Bigl{(}\sqrt{1-\text{\scriptsize$\frac{1}{4}$}\alpha^{2}}-\sqrt{1-\alpha^{2}}\Bigr{)}\,\chi_{*}$ (48)
instead of the result (34) of our scheme. Likewise, in their scheme, the decay rate of spontaneous emission from the excited state $\left|n=2,\,j=3/2\right\rangle$ to the groundstate $\left|n=1,\,j=1/2\right\rangle$ reads
$A=\frac{4\alpha}{3}\frac{m_{\text{FW}}^{3}\,c_{\text{FW}}^{4}}{h_{\text{FW}}^{3}}\left\langle\vec{r}_{if}\right\rangle^{2}\propto\frac{\chi_{*}^{3}.1^{4}}{1^{3}}\,\chi_{*}^{-2}\propto\chi_{*}$ (49)
instead of Eqs. (40)–(41) of our scheme.
Therefore, despite having an identical matter action, Eq. (14), our scheme and the Fujii–Wetterich scheme result in decisively different predictions for the behavior of the clock rate. Future technologies may be able to distinguish the two predictions, and hence the validity of our scheme versus that of the Fujii–Wetterich scheme.
Advantages of our mechanism—Our approach offers two distinct advantages over the Fujii–Wetterich treatment [7, 8, 9, 10, 11, 12]:
  1. 1.
    The Fujii–Wetterich scheme treats particle masses and charges at disparity: while masses are promoted to fields dependent on $\chi_{*}$, charges remain as fixed parameters (see Eq. (44)). Given that both masses and charges are intrinsic properties of particles, our scheme treats them on equal footing: particle masses and charges are parameters but not fields. (Note: in our scheme, $m$ and $e$ can ‘run’ in the RG flow, but they are independent of the background scalon $\chi$.)
  2. 2.
    The Fujii–Wetterich scheme only applies to massive particles, leaving massless particles unaffected. In contrast, the varying $c$ and $\hbar$ in our scheme impact all particles—massive and massless alike. Specifically, the variation in $c$ influences the propagation of (massless) photon in an expanding cosmic void, a feature absent in the Fujii–Wetterich scheme.
The anomalous $3/2$–exponent in our time scaling (27) is a novel discovery, leading to new physics with a unique set of previously unexplored phenomena and predictions. Specifically, it induces the variability of $c$ in spacetime, per $c_{*}:=l/\tau\propto\chi_{*}^{1/2}$, causing lightwaves traveling through the expanding cosmic void to undergo an additional refraction effect. This effect thence necessitates a reanalysis of the Pantheon Catalog of Type Ia supernovae, a task we will address in Section II.
II Phenomenology of VSL
This section focuses on our application of variable speed of light (VSL) for late-time cosmology, bypassing the need for dark energy (DE). The detailed work, including full technical elaboration, is presented in two companion papers, Refs. [14, 15].
I. VSL cosmography and cosmology—Our starting point is the generalized Friedmann–Lemaître–Robertson–Walker (FLRW) metric for the expanding cosmic void (with $\kappa=\{1,0,-1\}$)
$ds^{2}$ $=c^{2}\,dt^{2}-a^{2}(t)\,\left(\frac{dr^{2}}{1-\kappa\,r^{2}}+r^{2}\,d\Omega^{2}\right)$ (50)
$d\Omega^{2}$ $=d\theta^{2}+\sin^{2}\theta\,d\phi^{2}$ (51)
in which the speed of light $c$ is a function of the scalon field $\chi$ (per $c\propto\chi^{1/2}$ as specified in Eq. (21)). In the pure $\mathcal{R}^{2}$ gravity theory, Eq. (12), the scalon field is directly related to the Ricci scalar as $\mathcal{R}=2f_{0}\,\chi^{2}$, whereas for the metric (50) it can be shown that $\mathcal{R}\propto a^{-2}$. Consequently, in the cosmic void where the generalized FLRW metric is applicable, the following relation holds  [^12] This relation is a special case of a more generic form $c\propto a^{-\zeta}$ which Barrow heuristically proposed in [40]. In [15] we consider Barrow’s form in the context of power-law cosmology [41, 42].
$c\propto\chi^{1/2}\propto a^{-1/2}\,.$ (52)
The modified FLRW metric suitable for our VSL cosmology is thus expressed as
$ds^{2}=\frac{c_{0}^{2}}{a(t)}\,dt^{2}-a^{2}(t)\,\left(\frac{dr^{2}}{1-\kappa\,r^{2}}+r^{2}\,d\Omega^{2}\right)$ (53)
where the cosmic scale factor at our current time $t_{0}$ is set equal to 1, and $c_{0}$ is the speed of light measured in the void at $t_{0}$. It is important to note that several authors have previously applied a VSL cosmology to SnIa standard candles, while circumventing dark energy [43, 46, 44, 49, 48, 50, 47, 45, 51]. The consensus among these studies is that VSL alone could not adequately account for the SnIa data. However, these conclusions warrant reconsideration, as all previous analyses implicitly assumed the speed of light depends solely on cosmic time $t$. This assumption fails to hold for our VSL cosmology whereby $c$ varies in both space and time, making it location-dependent. This is because the scalon field $\chi$ can vary in spacetime, with its value in the intergalactic void (which is subject to cosmic expansion) differing from that in gravitationally bound galaxies (which host SnIa and are resistant to cosmic expansion). Hence, as $c\propto\chi^{1/2}$, $c$ is not only time-dependent but also location-dependent. A thorough treatment of this intricate issue is provided in our companion paper [14].
In the presence of VSL, the classic Lemaître redshift formula, $1+z=a^{-1}$, in standard cosmology is no longer applicable. Previous analyses of SnIa using VSL overlooked this crucial distinction and continued to use the classic formula, resulting in incorrect conclusions. The dependence of $c$ on the scalon field $\chi$ introduces two fundamental modifications to the right hand side of the Lemaître redshift relation, as follows:
  1. 1.
    The standard $a^{-1}$ term is replaced with $a^{-3/2}$, reflecting the $3/2$–exponent associated with the anisotropic time scaling, Eq. (27).
  2. 2.
    An additional correction accounts for the variation of the scalon field among galaxies as the function of redshift $z$. This is because galaxies hosting SnIa may possess different values for the scalon field due to their evolution in the expanding cosmic background. We model this physically motivated effect by introducing a monotonic function $F(z)$ that smoothly interpolates between two extremes: $F(z\rightarrow 0)=1$ and $F(z\rightarrow\infty)=F_{\infty}$ (with $F_{\infty}$ being an adjustable parameter). Specifically, we find the following formulation to be suitable
    $F(z)=1+(1-F_{\infty})\bigl{(}1-(1+z)^{-2}\bigr{)}^{2}\,.$ (54)
    By virtue of the length scaling (26), $F(z)$ thus signifies the evolution in the typical size of galaxies in response to the expansion of the comic void. This astronomical effect will have implications for the $H_{0}$ tension, which will be shown later in Point VIII.
Derivations of these modifications are detailed in our companion paper [14]. Here, we only quote the results. In our VSL cosmology, the modified Lemaître redshift formula reads
$1+z=a^{-3/2}\,F(z)\,.$ (55)
At very low $z$, with $a(t)=1-H_{0}\,d/c_{0}+\dots$ and $F(z)\approx 1+4(1-F_{\infty})z^{2}+\dots$, the relation (55) yields a modified Hubble law
$z\approx\frac{3}{2}H_{0}\,\frac{d_{L}}{c_{0}}\,.$ (56)
This significantly contrasts with the usual Hubble law $z\approx H_{0}\,d/c$ in standard cosmology. The multiplicative $3/2$–factor in Eq. (56) will have crucial implications in the estimation of $H_{0}$, as we shall see in Points III and V.
In a scale-invariant cosmology, the cosmic scale factor evolves as
$a=\left(\frac{t}{t_{0}}\right)^{2/3}$ (57)
reflecting the generic time scaling rule, $\tau\propto l^{3/2}$, derived in Eq. (27). Note that this evolution also occurs in the flat $\Lambda$CDM model in absence of dark energy ($\Omega_{\Lambda}=0$). Combining (55) and (57), we obtain the modified luminosity distance–redshift relation applicable for our VSL cosmology [14]
$\frac{d_{L}}{c_{0}}=\frac{1+z}{H_{0}\,F(z)}\,\ln\frac{(1+z)^{2/3}}{F(z)}\,.$ (58)
For comparison, this formula fundamentally differs from the canonical ones in the flat $\Lambda$CDM model ($\Omega_{M}+\Omega_{\Lambda}\!=\!1$)
$\frac{d_{L}}{c}=\frac{1+z}{H_{0}}\int_{0}^{z}\frac{dz^{\prime}}{\sqrt{\Omega_{M}(1+z^{\prime})^{3}+\Omega_{\Lambda}}}\,,$ (59)
and in the flat Einstein–de Sitter (EdS) universe (i.e. $\Omega_{M}=1,\,\Omega_{\Lambda}=0$)
$\frac{d_{L}}{c}=2\,\frac{1+z}{H_{0}}\left(1-\frac{1}{\sqrt{1+z}}\right)\,.$ (60)
II. Analyzing Pantheon Catalog of SnIa using VSL—The modified $d_{L}$-vs-$z$ relation (58), supplemented with (54), now stands ready for application to the Pantheon Catalog of SnIa. Our detailed analysis is provided in Ref. [14]. The Pantheon dataset, as provided in [52], is shown as open circles in the Hubble diagram in Fig. 1. It comprises $N=1,048$ data points given in terms of redshift and luminosity distance $\{z,\,d_{L}\}$, together with an error bar $\sigma_{L}$ of $d_{L}$, for $z$ spanning the range (0, 2.26)02.26(0,\,2.26)( 0 , 2.26 ). Fora given theoretical model, one would minimize the error
$\chi^{2}:=\frac{1}{N}\sum_{i=1}^{N}\biggl{(}\frac{d_{L}^{\text{model}}-d_{L}^{\text{Pantheon}}}{\sigma_{L}^{\text{Pantheon}}}\biggr{)}^{2}\,.$ (61)
Figure 1: Hubble diagram of SnIa in the Pantheon Catalog. Open circles: 1,048 data points with error bars, listed in Ref. [52]. In both panels, long-dashed line is $\Lambda$CDM Formula (59) with $H_{0}=70.2$, $\Omega_{M}=0.285$, $\Omega_{\Lambda}=0.715$; dotted line is flat CDM Formula (60) with $H_{0}=70.2$. Upper panel: Solid line is our VSL Formula (58) with $H_{0}=47.2$ and $F_{\infty}=0.93$. Lower panel: Solid line is our VSL Formula (58) with $F(z)\equiv 1\ \forall z$ with $H_{0}=44.4$.
Let us first review the canonical interpretation of SnIa data in standard cosmology. For very low $z$, both formulae (59) and (60) yield $\frac{d_{L}}{c}\approx\frac{z}{H_{0}}$. However, for positive $\Omega_{\Lambda}$, given the same $H_{0}$, the flat $\Lambda$CDM formula (59) produces a higher value for $d_{L}$ compared to the flat EdS formula (60). The crossover occurs at a redshift $z_{*}$ which approximately satisfies $\Omega_{M}(1+z_{*})^{3}\approx\Omega_{\Lambda}$. We fit Formula (59) to the Pantheon Catalog and obtain $H_{0}=70.2$, $\Omega_{M}=0.285$, $\Omega_{\Lambda}=0.715$ with an error $\chi_{\text{min}}^{2}=0.98824$. Hence, starting at $z\gtrsim z_{*}\approx 0.36$, the flat $\Lambda$CDM model begins to show an excess in $d_{L}$, meaning high-$z$ SnIa’s appear dimmer than the flat EdS model would predict. This behavior, manifest in the high-$z$ portion of the Hubble diagram (Fig. 1), has been interpreted as evidence for late-time cosmic accelerating expansion [58, 59] and for the existence of dark energy.
Our VSL formula (58) involves two adjustable parameters $H_{0}$ and $F_{\infty}$. Its best fit to the Pantheon Catalog yields $H_{0}=47.2$, $F_{\infty}=0.93$ with an error $\chi_{\text{min}}^{2}=0.98556$, which is highly competitive in quality to the $\Lambda$CDM fit. Note that in the upper panel of Fig. 1, the two curves—VSL and $\Lambda$CDM—are indistinguishable in the range $z\lesssim 2.26$ available for the Pantheon data  [^13] Light sources with even higher $z$, such as quasars [53], might help distinguish Formula (58) versus Formula (59)..Both curves manifest an excess in $d_{L}$ in the high-$z$ segment of the Hubble diagram. For completeness, we also show $F(z)$ and $F(a)$ as solid curves in Fig. 3.
Figure 2: Solid curves: $F(z)$ and $F(a)$. Dashed curves: $H_{0}(z)/H_{0}(z\text{=}0)$ and $H_{0}(a)/H_{0}(a=1)$.
Figure 3: Posterior distributions of $H_{0}$ (left panel) and $t_{0}$ (right panel), with 68% CL and 95% CL bands shown.
Of the two modifications to our modified Lemaître redshift formula $1+z=a^{-3/2}\,F(z)$ given in (55), the $a^{-3/2}$ term is expected to be of primary significance, while the $F(z)$ plays a secondary role. To test this intuition, we disable $F(z)$ by setting $F_{\infty}=1$ (making $F(z)\equiv 1\ \forall z\in\mathbb{R}$). A refit of (58) to the Pantheon data with just one parameter (i.e. $H_{0}$) yields $H_{0}=44.4$ with $\chi_{\text{min}}^{2}=1.25366$. Despite a substantial decrease in quality of fit, the $a^{-3/2}$ term alone still produces a reasonable fit, as shown in the lower panel of Fig. 1. Most importantly, this simplified fit continues to exhibit the excess in $d_{L}$ prominent in the high-$z$ segment of the Hubble diagram.
This indicates that our analysis practically—and parsimoniously—requires only one parameter $H_{0}$. The function $F(z)$ plays a role in resolving the $H_{0}$ tension, which will be explained in Point VIII below.
[…]
One pertinent question is the origin of this reduction in $H_{0}$. As we explained in Point I, for very low $z$, the Hubble law for our VSL cosmology is modified to $z\approx\frac{3}{2}H_{0}\,d/c_{0}$, as opposed to the standard Hubble law $z\approx H_{0}\,d/c$. Thus, for our VSL cosmology, the numerical value (empirically known to be $\sim 71$) for the slope of the $z$-vs-$d$ relation is not $H_{0}$ but $\frac{3}{2}H_{0}$, which promptly leads to $H_{0}\sim 47$. In conclusion, it is the $3/2-$exponent in the $a^{-3/2}$ term in our modified Lemaître redshift formula (55) that is responsible for the reduction in the $H_{0}$ value by a factor of $3/2$ compared to the canonical range.
Surprisingly, this reduced value of $H_{0}$ finds corroboration in a remarkable discover made by Blanchard et al in 2023 [54], when these authors analyzed the CMB power spectrum. We shall review their work below.
IV. Revitalizing Blanchard et al (BDRS)’s analysis of CMB power spectrum, bypassing DE—It is well established that the $\Lambda$CDM, with the primordial fluctuations spectrum $P(k)\propto k^{n}$, successfully accounts for Planck’s thermal power spectrum of the Cosmic Microwave Background (CMB) with parameters $n\approx 0.96$, $\Omega_{\Lambda}\approx 0.7$, $H_{0}\approx 67$. However, in 2003, Blanchard, Douspis, Rowan-Robinson, and Sarkar (BDRS) made a surprising discovery [54]. Instead of adhering to the standard spectrum $P(k)\propto k^{n}$, they considered two separate power-law forms for the low $k$ and high $k$ regimes, viz.
$P(k)=\begin{cases}A_{1}\,k^{n_{1}}&\text{for }k\leqslant k_{*}\\ A_{2}\,k^{n_{2}}&\text{for }k\geqslant k_{*}\end{cases}$ (62)
where $k_{*}$ is a break point, and continuity across $k_{*}$ is enforced by the condition $A_{1}\,k_{*}^{n_{1}}=A_{2}\,k_{*}^{n_{2}}$. Most importantly, they deliberately set $\Omega_{\Lambda}=0$, thereby avoiding the dark energy hypothesis. Remarkably, their model achieved an excellent fit to WMAP’s CMB data available at the time, with the optimal parameter values including $H_{0}=46$, $k_{*}=0.0096$ Mpc-1, $n_{1}=1.015$, $n_{2}=0.806$ [54]. Of these parameters, the most striking aspect was the low Hubble value $H_{0}=46$. A subsequent in-depth study by Hunt and Sarkar [55], utilizing a supergravity-based multiple inflation scenario, also successfully accounted for the CMB power spectrum without invoking dark energy, while requiring a similar value of $H_{0}\approx 43.5$. (Interestingly, Shanks in 2004 [56] also ventured that if $H_{0}\lesssim 50$ then a simpler, inflationary model with $\Omega_{baryon}=1$ might be allowed with no need for dark energy or cold dark matter.)
[…]
It is also important to note that the perceived low $H_{0}$ value $\sim 46$ eventually led BDRS to abandon their original findings in a follow-up study related to SDSS’s two-point correlation data of luminous red galaxies [57]. However, their abandonment may have been premature, since in the VSL framework the anomalous $3/2$–exponent in the anisotropic time scaling influences the observation of distant light sources along the time direction. Therefore, a reassessment of BDRS’s work of SDSS data, incorporating this effect, would be warranted. Additionally, future investigations should explore the impacts of VSL on phenomena such as lensing, Baryon Acoustic Oscillations (BAO), and the Etherington distance duality relation.
V. Resolving the age problem, bypassing DE—The flat $\Lambda$CDM model gives the age formula
$t_{0}^{\Lambda\text{CDM}}=\frac{2}{3\sqrt{\Omega_{\Lambda}}H_{0}}\,\text{arcsinh}\sqrt{\frac{\Omega_{\Lambda}}{\Omega_{M}}}\,.$ (63)
With $H_{0}=70.2$, $\Omega_{M}=0.285$, $\Omega_{\Lambda}=0.715$, it yields an age of $13.6$ Gy, an accepted figure in standard cosmology. If dark energy is absent ($\Omega_{\Lambda}=0,\,\Omega_{M}=1$), Eq. (63) simplifies to $t_{0}^{\text{EdS}}=2/(3H_{0})$ which then, with $H_{0}\sim 71$, would result in an age of $9.2$ Gy which would be too short to accommodate the existence of the oldest stars—a paradox commonly referred to as the age problem.
However, our VSL cosmology naturally resolves the age problem without requiring dark energy. Using the definition $H_{0}:=\left.\frac{1}{a}\frac{da}{dt}\right|_{t=t_{0}}$ and the evolution law (57), the age of the our VSL universe is
$t_{0}^{\text{VSL}}=\frac{2}{3H_{0}}\,.$ (64)
[…]
VI. Obtaining late-time acceleration via VSL, bypassing DEIt is an empirical fact that high-$z$ SnIa’s appear fainter than what would be predicted by the (matter-dominated) flat EdS model [58, 59]. This phenomenon is reflected in the distance modulus excess observed in the high-$z$ segment of the Hubble diagram in Fig. 1. Standard cosmology attributes this behavior to late-time cosmic acceleration, necessitating the introduction of a cosmological constant $\Lambda$ into the Friedmann equations, resulting in the $\Lambda$CDM model where $\Omega_{\Lambda}$ quantifies the amount of dark energy.
However, the analysis presented in Points I and II above demonstrates the efficacy of our VSL cosmology in accounting for the excess in $d_{L}$ without invoking dark energy. Its success can be quantitatively understood as follows. At high $z$, the luminosity distance formula of the flat EdS model (60) behaves as
$d_{L}\propto z\,.$ (65)
In contrast, the modified formula for our VSL cosmology (58) behaves as
$d_{L}\simeq z\,\ln z$ (66)
which gains an additional $\ln z$ term compared to (65). Consequently, according to our VSL Formula, $d_{L}$ grows faster and exhibits a steeper upward slope in the high-$z$ section in the Hubble diagram. Furthermore, this explanation can be strengthened through a qualitative account, as we will illustrate below.
VII. Physical intuition of late-time acceleration via VSL—Consider two supernovae, $A$ and $B$, located at distances $d_{A}$ and $d_{B}$ from the Earth, with $d_{B}=2\,d_{A}$. In standard cosmology, their redshift values $z_{A}$ and $z_{B}$ are related as $z_{A}\approx 2\,z_{A}$ (to first-order approximation). However, this relationship breaks down in our VSL cosmology. Since $c\propto a^{-1/2}$, light traveled faster in the distant past (when the cosmic factor $a\ll 1$) than in the more recent epoch (when $a\lesssim 1$). Therefore, photons emitted from supernova $B$ could cover twice the distance in less than double the time required by photons emitted from supernova $A$. Having spent less time in transit than predicted by standard cosmology, the $B$–photons experience less cosmic expansion than expected, resulting in a lower redshift such that:
$z_{B}<2\,z_{A}\ \ \text{for}\ \ d_{B}=2\,d_{A}\,.$ (67)
Conversely, consider a supernova $C$ with $z_{C}=2\,z_{A}$. For photons emitted from supernova $C$ to experience twice the redshift of the $A$–photons, they must travel a distance greater than twice that of the $A$–photons:
$d_{C}>2\,d_{A}\ \ \text{for}\ \ z_{C}=2\,z_{A}\,.$ (68)
This occurs because the $C$–photons traveled faster at the beginning of their journey toward Earth and must originate from a farther distance (thus appearing fainter than expected) to accumulate sufficient redshift. In either case, the VSL-based Hubble diagram exhibits a steeper upward slope in the high-$z$ section.
An alternative explanation for late-time acceleration is a decrease in the speed of light in the intergalactic void as the universe expands. In summary, VSL cosmology can account—qualitatively and quantitatively—for the distance modulus excess observed in high-$z$ SnIa without the need for dark energy. This approach overcomes the fine-tuning and coincidence problems associated with the concordance $\Lambda$CDM model. Furthermore, it bypasses the enigmatic nature of dark energy and aligns with BDRS’s 2003 groundbreaking analysis of the CMB power spectrum which also negates the dark energy hypothesis [54].
VIII. Resolving the $H_{0}$ tension via an astronomical origin—In recent years, interest in the $H_{0}$ tension has intensified [60]. One popular approach to address this issue involves treating $H_{0}$ as a “running” value, meaning that its value depends on the redshift of the data used to deduce it; e.g., see Ref. [61].
Our VSL cosmology naturally produces a “running” $H_{0}$ through the function $F(z)$. As mentioned in Point I, $F(z)$ represents the change in the typical size of galaxies as they evolve in an expanding intergalactic void (see also Ref. [14] for clarifications). From the modified formula (58), we can define the “running” value for $H_{0}$ as
$\frac{d_{L}}{c_{0}}=\frac{1+z}{H_{0}(z)}\,\ln(1+z)^{2/3}\,.$ (69)
By comparing this with Eq. (58) itself, we find
$H_{0}(z)=H_{0}\,F(z)\,\biggl{(}1-\frac{3}{2}\frac{\ln F(z)}{\ln(1+z)}\biggr{)}^{-1}\,.$ (70)
If $F(z)\equiv 1\ \forall z>0$, then $H_{0}(z)\equiv H_{0}$, a constant. Otherwise, as $z\rightarrow\infty$, $H_{0}(z)\rightarrow H_{0}F_{\infty}$; namely, at very high $z$, its “running” value is reduced by a factor $F_{\infty}$.
Using (54) and the value $F_{\infty}=0.93$ obtained from the fit in Point II, the computed “running” $H_{0}(z)$ is depicted as dashed curves in Fig. 3. At the time of recombination, the “running” value for $H_{0}$ would be $H_{0}\,F_{\infty}=47.2\times 0.93=43.9$, representing a $7\%$ reduction from its value estimated using low-redshift SnIa. The function $F(z)$ hence shed important light on the $H_{0}$ tension. Furthermore, in [55] Hunt and Sarkar obtained $H_{0}\approx 43.5$ from the CMB power spectrum using a supergravity-based multiple inflation scenario. Our result of $H_{0}(z\gg 1)\approx 43.9$ is thus in agreement with that of the Hunt–Sarkar team.
Importantly, the function $F(z)$ is a well-defined and physically motivated concept: it models the evolution of galaxies in an expanding intergalactic void. Consequently, the “running” $H_{0}$ in our approach is rooted in an astronomical origin, rather than a cosmic origin.
IX. An infinite cosmological horizon—In the generalized FLRW metric (50), the cosmological horizon at the recombination time $t_{rec}$ is given by
[…]
Since $c=c_{0}\,a^{-1/2}$ and $a=\left(t/t_{0}\right)^{2/3}$, we find
$l_{H}\left(t_{rec}\right)$ $=c_{0}\,t_{0}\,\left(\frac{t_{rec}}{t_{0}}\right)^{2/3}\int_{0}^{t_{rec}}\frac{dt^{\prime}}{t^{\prime}}=\infty\,.$ (72)
The cosmological horizon is thus (logarithmically) divergent, suggesting that the entire universe was causally connected at the recombination event. This result dovetails with the VSL proposals previously made by Moffat [21] and Albrecht and Magueijo [22] which sought to explain the near isotropy of the CMB temperatures—an empirical fact commonly known as the horizon problem.
III Conclusions
The action given in Eqs. (6), (7), and (12) describes a generally covariant theory of gravitation and matter, which also possesses dilatation symmetry. A notable feature is the non-minimal coupling of matter with gravity, expressed through the coupling of the Higgs field $\Phi$ with a background ‘scalon’ field $\chi$ in the form $\chi^{2}\,\Phi^{2}$. Crucially, the (dimensionful) quantum of action $\hbar$ and speed of light $c$ are absent in this adimensional action. Previously, Fujii and Wetterich utilized this action to generate variable particle masses and (Newton) gravitational ‘constant’ [8, 9, 10, 11, 12]. In Section I of this paper, we developed an alternative mechanism to construct (variable) $\hbar$ and $c$ from the background scalon field $\chi$. Our approach can be generalized when the matter Lagrangian (7) is extended to cover the Glashow–Weinberg–Salam (GWS) model of the $SU(3)\times SU(2)\times U(1)$ gauge group [13].
In our framework, each open set vicinity of any given point $x^{*}$ on the spacetime manifold is equipped with a replica of the GWS action for matter fields (fermions, gauge vector bosons, and Higgs scalar bosons), operating with its own effective values of $\hbar_{*}$ and $c_{*}$. Lorentz symmetry holds locally within each open set. These GWS replicas are isomorphic (i.e. structurally identical), differing only in the values of $\hbar_{*}$ and $c_{*}$, which are determined by the prevailing value of the background scalon field $\chi(x^{*})$ per $\hbar_{*}=\chi^{-1/2}(x^{*})$ and $c_{*}=\chi^{1/2}(x^{*})$. Crucially, within each open set, the length and time scales of physical processes are governed by the scalon field as $l\propto\chi^{-1}(x^{*})$ and $\tau\propto\chi^{-3/2}(x^{*})$, resulting in an anomalous anisotropic relation between the rate of clocks and the length of rods, $\tau^{-1}\propto l^{-3/2}$. This novel relation enables us to predict a new “time dilation effect of the Third kind”, with further details presented in Ref. [16].
Our construction hence fulfills Einstein’s original—albeit long overlooked—vision in incorporating the local validity of Lorentz symmetry, the variation of speed of light in curved spacetimes, and the geometric nature of gravity [18].
The anisotropic relation $\tau^{-1}\propto l^{-3/2}$ mentioned above gives rise to new physics and phenomenology. Section II of our paper presented key findings on the phenomenological consequences of our mechanism in the context of cosmology, particularly addressing late-time cosmic acceleration via the variation in $c$ while eliminating the need for dark energy. Remarkably, our results revive a crucial—but previously overlooked—discovery made by Blanchard, Douspis, Rowan-Robinson, and Sarkar in 2003 which circumvented the requirement for dark energy in their analysis of the CMB power spectrum [54].
Acknowledgements.
I wish to thank Clifford Burgess for his instrumental feedback, which greatly sharpened the exposition of this paper. I also thank Anne-Christine Davis and Sergey Sibiryakov for their valuable insights. The constructive comments from Robert Mann, Tiberiu Harko, Tessa Baker, John Moffat, Niayesh Afshordi, Alexander Kamenshchik, Bertrand Chauvineau, Radosław Kycia, Péter Ván, Lorenzo Gavassino, and Gregory Horndeski are gratefully acknowledged.
Appendix:  On Planck’s and Hartree’s unit systems
In our mechanism, the electron mass $m$ and charge $e$ are dimensionless quantities derived from the parameters $f$, $\mu$, $\lambda$, and $\alpha$ via Eq. (19). The scalon field $\chi$ is the single entity that carries a scale, which then serves as the only unit available. Consequently, all physical quantities can be expressed exclusively in terms of the magnitude of $\chi$. For instance, $\hbar$ and $c$ have the units of $\chi^{-1/2}$ and $\chi^{1/2}$, respectively, whereas $G$ is dimensionless. The rest mass energy $E=mc^{2}$ has the unit of $\chi$, owing to $m$ being dimensionless and $c\propto\chi^{1/2}$. Similarly, the Coulomb potential of a point charge $e$, given by $e^{2}/r$, also carries the unit of $\chi$, by virtue of $e$ being dimensionless and $r\propto\chi^{-1}$. In general, a theory with dilatation symmetry would require only one ‘unit’ (associated with the scalon field $\chi$), rendering the canonical three-unit system (mass $M$, length $L$, time $T$) unnecessary.
Nevertheless, to distinguish between mass and charge, which represent two distinct aspects of the electron (viz., inertia and $U(1)$ gauge coupling strength), we introduce two ‘labels’ $m_{0}$ and $e_{0}$ into their definitions, Eq. (19)
$m:=\frac{f\mu}{\sqrt{\lambda}}\,m_{0};\ \ \ e:=\sqrt{\alpha}\,e_{0}\,.$ (73)
This redefinition leads to
$\hbar_{*}:=e_{0}\sqrt{m_{0}}\,\chi_{*}^{-1/2};\ \ c_{*}:=\frac{e_{0}}{\sqrt{m_{0}}}\,\chi_{*}^{1/2};\ \ G:=\frac{e_{0}^{2}}{m_{0}^{2}}\,.$ (74)
Our new unit system based on the trio $\{e_{0},\,m_{0},\,\chi_{*}\}$ deviates from the Planck $cGh$ unit system, which relies on the trio of constants $\{c,\,G,\,\hbar\}$. Our new unit system is akin to the unit system introduced by Hartree in 1928 [62], based on the electron charge $e$, mass $m$, and the Bohr radius $a_{B}$ (the typical size of the hydrogen atom in its ground state). In the Hartree system:
$\hbar$ $=e\sqrt{m\,a_{B}};\ \ \ c=\frac{e}{\alpha\sqrt{ma_{B}}}\,.$ (75)