### v3

# A mechanism to generate varying speed of light via Higgs-dilaton coupling:Theory and cosmological applications

Abstract
We probe into a class of scale-invariant actions, which allow the Higgs field $\Phi$ to interact with a dilaton field $\chi$ of the background spacetime through the term $\chi^{2}\,\Phi^{\dagger}\Phi$. Upon spontaneous gauge symmetry breaking, the vacuum expectation value (VEV) of the Higgs field becomes proportional to $\chi$. Although this linkage is traditionally employed to make the Planck mass and particle masses dependent on $\chi$, we present an alternative mechanism: the Higgs VEV will be used to construct Planck’s quantum of action $\hbar$ and speed of light $c$. Specifically, each open set vicinity of a given point $x^{*}$ on the spacetime manifold is equipped with a replica of the Glashow–Weinberg–Salam action operating with its own effective values of $\hbar_{*}$ and $c_{*}$ per $\hbar_{*}\propto\chi^{-1/2}(x^{*})$ and $c_{*}\propto\chi^{1/2}(x^{*})$, causing these “fundamental constants” to vary alongside the dynamical field $\chi$. Moreover, in each open set around $x^{*}$, the prevailing value $\chi(x^{*})$ determines the length and time scales for physical processes occurring in this region as $l\propto\chi^{-1}(x^{*})$ and $\tau\propto\chi^{-3/2}(x^{*})$. This leads to an anisotropic relation $\tau^{-1}\propto l^{-3/2}$ between the rate of clocks and the length of rods, resulting in a distinct set of novel physical phenomena. For late-time cosmology, the variation of $c$ along the trajectory of light waves from distant supernovae towards the Earth-based observer necessitates modifications to the Lemaître redshift formula, the Hubble law, and the luminosity distance–redshift relation. These modifications are capable of: (1) Accounting for the Pantheon Catalog of Type Ia supernovae through a declining speed of light in an expanding Einstein–de Sitter universe, thus avoiding the need for dark energy; (2) Revitalizing Blanchard–Douspis–Rowan-Robinson–Sarkar’s CMB power spectrum analysis that bypassed dark energy [A&A 412, 35 (2003)]; and (3) Resolving the $H_{0}$ tension without requiring a dynamical dark energy component.
I Motivation
In light of several outstanding issues and tensions in cosmology, growing interest has emerged in promoting the parameters of physical models to be variable scalar fields in spacetime [1, 4, 2, 3]. For example, the cosmological constant $\Lambda$ is equipped with a kinetic term that allows it to evolve and ‘relax’ to its (small) value in the current epoch, thus potentially offering a relief to the fine-tuning and coincidence problems in late-time cosmology [5, 6].
Against this backdrop, the accustomed Planck $cGh$ unit system (1899-1900) solidifies the role of the speed light $c$, the (Newton) gravitational constant $G$, and the quantum of action $\hbar$ as “fundamental constants” which encompass the realms of special relativity, gravitation, and quantum mechanics, respectively [7, 8, 9]. The trio $\{c,\,G,\,\hbar\}$ are viewed as fixed cornerstones, convertible to the three “fundamental units”—the Planck mass, the Planck length, and the Planck time, defined as
$M_{P}:=\sqrt{\frac{\hbar c}{G}};\ \ l_{P}:=\sqrt{\frac{\hbar G}{c^{3}}};\ \ \tau_{P}:=\sqrt{\frac{\hbar G}{c^{5}}}$ (1)
These units span the basis for describing all physical phenomena known to date.
This orthodox view has faced challenges, however. The most well-known example is the concept of variable $G$, inspired by Dirac in 1937 [10], and culminating in the generally covariant Brans-Dicke (BD) theory of gravity in 1961 [11]. In this theory, a dynamical scalar degree of freedom $\chi$ is introduced alongside the metric tensor $g_{\mu\nu}$ as
$\int d^{4}x\,\frac{\sqrt{-g}}{16\pi}\left[\chi^{2}\,\mathcal{R}-4\omega\,\nabla_{\mu}\chi\nabla^{\mu}\chi\right]$ (2)
where $\mathcal{R}$ is the Ricci scalar, and $\omega$ a dimensionless (BD) parameter. The BD theory has grown into a family of scalar–tensor theories, a very popular theme of research nowadays. The field $\chi$, often referred to as a ‘dilaton’, also naturally arises in string theory and other contexts [12].
Comparing Eq. (2) with the Einstein–Hilbert (EH) action
$\mathcal{S}_{\text{EH}}=\int d^{4}x\,\frac{\sqrt{-g}}{16\pi}\frac{c^{3}}{\hbar\,G}\,\mathcal{R}$ (3)
where the speed of light $c$ and quantum of action $\hbar$ are explicitly restored, the (Newton) gravitational ‘constant’ in BD theory becomes a scalar field via the following formal identification
$\chi^{2}:=\frac{c^{3}}{\hbar\,G}$ (4)
Under the canonical assumption that $c$ and $\hbar$ do not depend on $\chi$, Eqs. (1) and (4) lead to the relationships
$G=\frac{c^{3}}{\hbar}\chi^{-2};\ \ \ M_{P}=\frac{\hbar}{c}\chi;\ \ \ l_{P}=\chi^{-1};\ \ \ \tau_{P}=\frac{1}{c}\chi^{-1}$ (5)
The dilaton field $\chi$ hence directly determines the Planck length.
It is imperative to stress that, in BD theory, the field $\chi$ itself serves as a (dynamical) length scale that is used to construct $G$, a dimensionful quantity. However, the identity expressed in Eq. (4) permits an alternative mechanism: The field $\chi$, a quantity equal to $\left(c^{3}/(\hbar G)\right)^{1/2}$, can also be used to construct $c$ and $\hbar$, rather than $G$. The key objective of our paper is to establish the relations of $\hbar$ and $c$ with $\chi$ via the matter sector.
In the original BD theory, the matter sector coupled minimally to the gravitational sector to preserve Einstein’s Equivalence Principle [11]. However, around the same time, Utiyama and DeWitt showed that even if one starts with a minimal coupling action between matter and gravity, radiative corrections would generate non-minimal coupling between them [13]. Since the 1980s, efforts to introduce non-minimal coupling directly at tree level have been actively pursued in the literature. For example, in [14, 15, 16, 17, 18] Wetterich embedded the BD action in a broader framework that projects a dilatation symmetry, with the non-minimal coupling in the form  [^1] Wetterich referred the field $\chi$ a ‘cosmon’ [14]. In place of Eq. (6), he considered a more general form $V(\chi,\Phi)=\Phi^{4}\,w\left(\chi^{2}/\Phi^{2}\right)$ where $w$ is a function of the ratio $x:=\chi^{2}/\Phi^{2}$. If $w(x)$ adopts an affine form $\lambda-\mu^{2}\,x$, then Eq. (6) would ensue.
$V(\chi,\Phi)=\lambda\,\Phi^{4}-\mu^{2}\,\chi^{2}\,\Phi^{2}$ (6)
where $\lambda\in\mathbb{R^{+}}$ and $\mu\in\mathbb{R}$ are dimensionless parameters. In [19], a similar form was also considered in the context of Higgs–dilaton cosmology. Preceding to this, in [20], in place of $\chi^{2}$, Fujii let the Higgs doublet couple with the Ricci scalar in the form $\mathcal{R}\,\Phi^{\dagger}\Phi$.
In the presence of Higgs–dilaton coupling, when the electroweak gauge symmetry of the action undergoes spontaneous breaking, the field $\chi$ influences the vacuum expectation value (VEV) of the Higgs doublet $\Phi$, which in turn affects the mass of the fermions, the gauge vector bosons, and the Higgs bosons in the Glashow–Weinberg–Salam (GWS) model. Specifically, Fujii and Wetterich allowed the particle mass $m$ to scale proportional to $\chi$, thereby maintaining a constant ratio $m/M_{P}$ at classical level. We shall refer to their practice collectively as the Fujii–Wetterich (FW) scheme. Moreover, the length scale $l$ and the time scale $\tau$ of any given physical process also become dependent on $\chi$ as $l\propto\chi^{-1}$ and $\tau\propto\chi^{-1}$, in exact proportion to $l_{P}$ and $\tau_{P}$ respectively. Importantly, the field $\chi$ affects not only the Planck time per Eq. (5) but also the rate of physical clocks (viz. the revolution rate of a mechanical clock, the oscillation rate of an atomic clock, or the decay rate of an unstable quantum system) which are made of matter governed by the GWS model. Quantitatively, the FW scheme thus predicts that the evolution rate $\tau^{-1}$ of physical processes is proportional to $\chi$.  [^2] We find it illuminating to quote Fujii [20]: “… the time and length in the microscopic unit frame are measured in units of $m^{-1}(t)$, in agreement with the physical situation that the time scale of atomic clocks, for example, is provided by the atomic levels which are determined by the Rydberg constant $(me^{4})^{-1}$ and Wetterich [16]: “The clock provided by the Hubble expansion in the standard description is now replaced by a clock associated to the increasing value of $\chi$.
As stated earlier, our approach constitutes a major departure from the FW scheme: Instead of imposing that $\hbar$ and $c$ be constant, we require the charge and inertial mass of particles to be independent of $\chi$. This unambiguously leads to the relations $c\propto\chi^{1/2}$ and $\hbar\propto\chi^{-1/2}$, while $G$ remains constant. Furthermore, via the time evolution of quantum states $i\hbar\frac{\partial}{\partial t}\left|\psi\right\rangle=\hat{H}\left|\psi\right\rangle$, the dependency $\hbar\propto\chi^{-1/2}$, along with $\hat{H}\propto\chi$, causes the evolution rate of physical processes to scale as $\tau^{-1}\propto\chi^{3/2}$, in decisive distinction from the FW scheme (which posits $\tau^{-1}\propto\chi$).
The anisotropic time scaling enabled in our approach, characterized by its anomalous $3/2$–exponent, leads to new physics with a distinct set of phenomenology andpredictions. We will discuss these significant implications in this paper, with further details developed in Ref. [21].
In brief:  History of variable c and $\hbar$
Prior to the development of Brans-Dicke theory, Dicke also briefly considered variability in the speed of light $c$ in 1957 [22] although he did not pursue it further. A relatively little-known fact is that the concept of variable speed of light (VSL) was initially explored by Einstein in 1911, published in three obscure papers (originally in German) [23, 24, 25]  [^3] It seems that Dicke in 1957 [22] was not aware of Einstein’s 1911 VSL proposal [23]. during his quest for a generally covariant theory of gravity. In [24, 25] Einstein emphasized that the Michelson–Morley experimental results and Lorentz symmetry are meant to hold only locally; while $c$ is invariant with respect to local boosts, its value needs not be universal. Einstein envisioned the possibility that the gravitational potential affected the clock rate, hence generating a varying $c$ in spacetime. Such a variation in $c$ could then produce a refraction effect on light rays around the Sun’s disc as a result of Huygens’s principle [23]. The later successes of General Relativity (GR) quickly overshadowed Einstein’s 1911 paper on VSL, leading him to neglect the idea, however. In the 1990s, the VSL concept was revived independently by Moffat and by Albrecht and Magueijo to tackle issues in early-time cosmology, such as the horizon paradox [26, 27]. Since then, several researchers actively explore various aspects of VSL [29, 30, 31, 28, 49, 52, 51, 50, 39, 40, 68, 53, 54, 66, 67, 82, 96, 86, 63, 62, 61, 80, 55, 32, 77, 34, 88, 92, 70, 89, 65, 47, 36, 85, 93, 90, 94, 74, 81, 83, 87, 57, 56, 71, 91, 95, 58, 72, 69, 73, 97, 48, 46, 43, 84, 79, 35, 42, 41, 38, 37, 76, 75, 44, 60, 59, 78, 98, 99, 100, 101, 45, 33, 64].
The possibility of a varying quantum of action $\hbar$ has been much less explored, with a rare exception of [102]. One important empirical guidance is that, although the fine-structure constant $\alpha:=e^{2}/(\hbar c)$ is known to ‘run’ in the renormalization group (RG) flow of quantum field theory, there is no clear experimental evidence supporting $\alpha$ varying in spacetime  [^4] In [103, 104] Webb et al. reported observational evidence of a varying $\alpha$ with respect to the redshift, but this result has been in contention; e.g., see a recent paper [105]. In addition, it violates the tight theoretical bound given in [106].. If $\alpha$ is to remain independent of the field $\chi$ while $c$ is allowed to vary, it would require a co-variation of $e$, $\hbar$, and $c$. One such scenario, as briefly mentioned earlier, posits $\hbar\propto\chi^{-1/2}$ and $c\propto\chi^{1/2}$, which would insulate $\alpha$ from depending on $\chi$.
II Higgs–dilaton coupling
Consider the full action in a 4-dimentional spacetime:
$\mathcal{S}$ $=\int d^{4}x\,\sqrt{-g}\,\mathcal{L}_{\text{mat}}+\int d^{4}x\,\frac{\sqrt{-g}}{16\pi}\,\mathcal{L}_{\text{grav}}$ (7)
Let us first focus on the matter sector. As a prototype, we consider the following Lagrangian, representing a massless spinor field $\psi$ coupled with a massless $U(1)$ gauge field $A_{\mu}$, and a charge-neutral Higgs singlet $\Phi$ coupled with a dilaton $\chi$:
$\mathcal{L}_{\text{mat}}$ $=i\,\bar{\psi}\gamma^{\mu}(\nabla_{\mu}-i\sqrt{\alpha}A_{\mu})\psi+f\,\bar{\psi}\psi\,\Phi$
$+\frac{\mu^{2}}{2}\,\chi^{2}\,\Phi^{2}-\frac{\lambda}{4}\,\Phi^{4}+g^{\mu\nu}\partial_{\mu}\Phi\,\partial_{\nu}\Phi-\frac{1}{4}\,F_{\mu\nu}F^{\mu\nu}$ (8)
All parameters $\alpha$, $f$, $\mu$, and $\lambda$ are dimensionless. It is important to remark that the adimensional action of the matter sector, $\int d^{4}x\sqrt{-g}\,\mathcal{L}_{\text{mat}}$, does not involve the (dimensionful) quantum of action $\hbar$ and speed of light $c$ at its outset. The length dimensions of $\{\psi,\,A_{\mu},\,\Phi$} are $\{-3/2,-1,-1\}$ respectively. The gamma matrices satisfy $\gamma^{\mu}\gamma^{\nu}+\gamma^{\nu}\gamma^{\mu}=2g^{\mu\nu}$, and the spacetime covariant derivative $\nabla_{\mu}$ acts on the spinor via vierbein and spin connection.
The Lagrangian $\mathcal{L}_{\text{mat}}$ in Eq. (8) is invariant under the local $U(1)$ gauge transformation:
$\psi(x)$ $\rightarrow e^{i\sqrt{\alpha}\sigma(x)}\,\psi(x)$ (9)
$A_{\mu}(x)$ $\rightarrow A_{\mu}(x)-\partial_{\mu}\sigma(x)$ (10)
where $\sigma(x)$ is an arbitrary scalar function. In addition, $\mathcal{L}_{\text{mat}}$ is also invariant under the discrete $\mathbb{Z}_{2}$ symmetry of the Higgs singlet
$\Phi\leftrightarrow-\Phi$ (11)
Regarding the gravitation Lagrangian $\mathcal{L}_{\text{grav}}$, most extended theories of gravitation naturally host a scalar degree of freedom [107, 108, 109, 12, 110, 111]. Several possibilities exist, one being a massive Brans-Dicke action
$\mathcal{L}_{\text{grav}}=\chi^{2}\,\mathcal{R}-4\omega\,\nabla^{\mu}\chi\nabla_{\mu}\chi-V(\chi)$ (12)
When $\omega=0$, $\mathcal{L}_{\text{grav}}$ in Eq. (12) belongs to the $f(\mathcal{R})$ family of gravity [112, 113, 114]. It can also be embedded within a broader family of scale invariant gravity, such as those considered in [115, 116, 117] or “agravity” which includes the Weyl term, $\mathcal{C}^{\mu\nu\lambda\rho}\,\mathcal{C}_{\mu\nu\lambda\rho}$ [119, 118, 120]. Alternatively, it can represent Horndeski gravity, the most general theory in four dimensions constructed out of the metric tensor and a scalar field that leads to second-order equations of motion [121]. Another viable option would be Lyra geometry [122], where the inherent scale function of this geometry can play the role of the dilaton field $\chi$. However, the exact details of the gravitational sector are not the focus of our paper. Our work concerns the matter sector, and we require only the existence of a ‘dilaton’ singlet $\chi$ that couples with the Higgs field $\Phi$ as described in Eq. (8).
It has been established in [123, 115] that a scale-invariant action, such as the one described in Eqs. (7), (8), and (12), can evade observational constraints on the fifth force. Among the several possible choices for $\mathcal{L}_{\text{grav}}$, the pure $\mathcal{R}^{2}$ theory $\mathcal{L}_{\text{R2}}=\chi^{2}\,\mathcal{R}-f_{0}\,\chi^{4}$ is a promising candidate [125, 126, 127, 124]. At the classical level, the dilaton and the Ricci scalar are one-to-one related as $\mathcal{R}=2f_{0}\,\chi^{2}$. The Newtonian limit has been established [128, 124]. The theory is known to be Ostrogradsky stable and free of ghosts [124], and properties of its graviton propagators have also been investigated [129, 125, 124].
III A new mechanism to generate variable $\hbar$ and c
Assuming that $\chi$ is a slowly varying background field, at a given point $x^{*}$ on the manifold, its value is $\chi_{*}:=\chi(x^{*})$. In the open set vicinity of the point $x^{*}$, the terms $\frac{\mu^{2}}{2}\,\chi_{*}^{2}\,\Phi^{2}-\frac{\lambda}{4}\,\Phi^{4}$ in Eq. (8), evaluated for the prevailing $\chi_{*}$, induce a spontaneous breaking of the discrete $\mathbb{Z}_{2}$ symmetry of the Higgs singlet, given by Eq. (11)  [^5] Note: The $U(1)$ gauge symmetry described by Eqs. (9)–(10) remains unbroken in this process..This process, traditionally known as the Higgs mechanism [130, 132, 133, 131], results in a non-zero vacuum expectation value (VEV) of the Higgs field
$\left\langle\Phi(x^{*})\right\rangle=\frac{\mu}{\sqrt{\lambda}}\chi_{*}$ (13)
which is proportional to $\chi_{*}$. In a local reference frame tangent to the manifold at the point $x^{*}$ (i.e, an inertial frame at $x^{*}$), the action for the matter sector (excluding the Higgs excitation above the vacuum) becomes  [^6] If we adopt the analysis presented in Ref. [134], the Higgs VEV $v(x):=\left\langle\Phi(x)\right\rangle$ would obey the equation $-2\square v+\mu^{2}\chi^{2}v-\lambda v^{3}=0$. However, we treat $v$ as a slowly varying background field; under this condition, the projection of the matter sector onto the local inertial frame, viz. Eq. (14), is justified.
$\int d^{4}x\,\biggl{[}i\bar{\psi}\gamma^{\mu}\partial_{\mu}\psi+\sqrt{\alpha}\,\bar{\psi}\gamma^{\mu}A_{\mu}\psi+\frac{f\mu}{\sqrt{\lambda}}\,\chi_{*}\,\bar{\psi}\psi-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}\biggr{]}$ (14)
Varying the spinor field and the gauge vector field, we obtain two equations of motion
$\biggl{(}i\gamma^{\mu}\partial_{\mu}+\sqrt{\alpha}\,\gamma^{\mu}A_{\mu}+\frac{f\mu}{\sqrt{\lambda}}\,\chi_{*}\biggr{)}\,\psi$ $=0$ (15)
and
$\partial_{\nu}F^{\nu\mu}=j^{\mu}:=\sqrt{\alpha}\,\bar{\psi}\gamma^{\mu}\psi$ (16)
The first equation resembles the Dirac equation for the ‘electron’ $\psi$, coupled with a $U(1)$ gauge vector $A_{\mu}$
$\biggl{(}i\gamma^{\mu}\partial_{\mu}+\frac{e}{\sqrt{\hbar c}}\,\gamma^{\mu}A_{\mu}+m\,\frac{c}{\hbar}\biggr{)}\,\psi=0$ (17)
The second equation resembles the Maxwell equation for $A_{\mu}$ sourced by a $U(1)$–charged current $j^{\mu}$
$\partial_{\nu}F^{\nu\mu}=j^{\mu}:=\frac{e}{\sqrt{\hbar c}}\,\bar{\psi}\gamma^{\mu}\psi$ (18)
In Eq. (17), the parameters $e$ and $m$ represent the $U(1)$ gauge charge and inertial mass of the electron, respectively. As they are intrinsic properties of the electron, we then require them to be parameters rather than fields, namely, they are independent of $\chi_{*}$. (Note: We must emphasize that $e$ and $m$ are not constants, but they can ‘run’ in the renormalization group flow when radiative corrections involving $\psi$ and $A_{\mu}$ are included.) By comparing Eqs. (15)–(16) against Eqs. (17)–(18), we can identify
$m:=\frac{f\mu}{\sqrt{\lambda}};\ \ \ e:=\sqrt{\alpha}$ (19)
These identities then lead to
$\hbar\,c:=1;\ \ \ \frac{c}{\hbar}:=\chi_{*}$ (20)
which unambiguously yield the following relations
$\hbar_{*}:=\chi_{*}^{-1/2};\ \ \ c_{*}:=\chi_{*}^{1/2}$ (21)
Here, the subscript * in $\hbar_{*}$ and $c_{*}$ signifies the dependence of $\hbar$ and $c$ as functions of $\chi$. It should be noted that Newton’s constant in independent of $\chi_{*}$ because $G:=\frac{c_{*}^{3}}{\hbar_{*}\chi_{*}^{2}}=1$.
Conversely, in the tangent—i.e., inertial—frame at the point $x^{*}$, the action in Eq. (14) can be recast as an action of Quantum Electrodynamics (QED) with a Lagrangian expressed in terms of $\hbar_{*}$, $c_{*}$, $m$, and $e$, viz.
$\mathcal{L}_{\text{QED}}=i\bar{\psi}\gamma^{\mu}\partial_{\mu}\psi+\frac{e}{\sqrt{\hbar_{*}c_{*}}}\bar{\psi}\gamma^{\mu}A_{\mu}\psi+m\frac{c_{*}}{\hbar_{*}}\bar{\psi}\psi-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}$ (22)
Therefore, each open set enclosing a given point $x^{*}$ on the manifold is equipped with a replica of the QED action (22), operating with an effective Planck constant $h_{*}$ and an effective speed of light $c_{*}$. Both of these effective parameters are determined by the prevailing value $\chi_{*}$ of the background dilaton field, per Eq. (21). As a component of the gravitational sector, $\chi_{*}$ can vary across the manifold, leading to corresponding variations in $\hbar_{*}$ and $c_{*}$ throughout the manifold.
Within each open set, the effective speed of light $c_{*}$ governs the propagation of the (massless  [^7] Note: Only the discrete $\mathbb{Z}_{2}$ symmetry (11) of the Higgs singlet is broken. The local $U(1)$ gauge symmetry in Eqs. (9) and (10) remains unbroken, leaving the gauge vector boson $A_{\mu}$ massless.) gauge vector field $A_{\mu}$, whereas the effective Planck constant $\hbar_{*}$ regulates the quantization of the fields $\psi$, $A_{\mu}$, and $\Phi$. We should note that our mechanism for generating variable $\hbar$ and $c$ is readily applicable to the generalization of $\mathcal{L}_{\text{mat}}$ in Eq. (8) to the full GWS model of particle physics, where the gauge group is enlarged to $SU(3)\times SU(2)\times U(1)$. This extension is a relatively straightforward exercise. A simplified version of our mechanism is also provided in Ref. [135] where we allowed the dilaton to interact directly with the fermion field in the form $\chi\,\bar{\psi}\psi$. We obtained results that are essentially identical to those presented above.
The role of $\hbar_{*}$ is instrumental: from Eq. (15) (and using $\hbar_{*}c_{*}=1$ per Eq. (21)), it can be shown that the time evolution of the electron wavefunction is given by
$i\hbar_{*}\frac{\partial}{\partial t}\psi(t)$ $=\hat{H}(t)\,\psi(t)$ (23)
$\hat{H}(t)$ $:=-i\alpha^{k}\frac{\partial}{\partial x^{k}}+\sqrt{\alpha}\left(\alpha^{k}A_{k}-\beta A_{0}\right)+\frac{f\mu}{\sqrt{\lambda}}\,\chi_{*}\,\beta$ (24)
Here, the matrices are $\alpha^{k}:=\gamma^{0}\gamma^{k}$ and $\beta:=\gamma^{0}$. The time evolution of $\psi(t)$ within the open set enclosing $x^{*}$ thus depends on the behavior of $\hbar_{*}$ with regard to $\chi_{*}$. This crucial connection between $\hbar_{*}$ and the evolution rate of quantum states leads to a concrete prediction, which we shall elaborate in the subsequent sections.
IV A prediction:   Anisotropic scaling in the clock rate
The effective QED action given in Eq. (14) possesses a dilatation symmetry; namely, it remains invariant under the transformation
$dx^{\mu}\rightarrow\chi_{*}\,dx^{\mu};\ \psi\rightarrow\chi_{*}^{-3/2}\psi;\ A_{\mu}\rightarrow\chi_{*}^{-1}A_{\mu}$ (25)
The prevailing value of the dilaton field $\chi$ at the point $x^{*}$ thus sets the length scale for a given physical process within the open set surrounding $x^{*}$, i.e.
$l\propto\chi_{*}^{-1}$ (26)
justifying the term “dilaton” for $\chi$. However, due to the dependence of $c$ on $\chi$ as given in Eq. (21), the timescale for the physical process in the open set exhibits an anisotropic behavior
$\tau:=\frac{\chi_{*}^{-1}}{c_{*}}\propto\chi_{*}^{-3/2}$ (27)
This behavior can also be understood through the time evolution operator (23)–(24). Since $dx^{k}\propto\chi_{*}^{-1}$ and $A_{\mu}\propto\chi_{*}$, the Hamiltonian in (24) scales as $\chi_{*}$. The rate of evolution is thus
$\tau\simeq\frac{\hbar_{*}}{\hat{H}}\propto\chi_{*}^{-3/2}$ (28)
which is compatible with the anomalous time scaling found in Eq. (27).
To illustrate the behavior of the length and time scales with respect to $\chi_{*}$, we will consider the Hydrogen atom as an example. The Bohr radius is
$a_{B}$ $=\frac{\hbar_{*}}{\alpha\,mc_{*}}=\frac{\sqrt{\lambda}}{\alpha f\mu}\,\chi_{*}^{-1}\propto\chi_{*}^{-1}$ (29)
which is consistent with Eq. (26). The energy level of (relativistic) electron in a quantum state $\left|n,\,j\right\rangle$ of a hydrogen atom is a well-established result [136]
$E_{n}^{j}$ $=\mathcal{N}_{n}^{j}\,mc_{*}^{2}=\mathcal{N}_{n}^{j}\,\frac{f\mu}{\sqrt{\lambda}}\,\chi_{*}$ (30)
in which $\mathcal{N}_{n}^{j}\!:=\!{\scriptstyle{\scriptstyle\left(1+\alpha^{2}\left(n-j-\frac{1}{2}+\sqrt{\left(j+\frac{1}{2}\right)^{2}-\alpha^{2}}\right)^{-2}\right)^{-1/2}}}$ with $n=1,2,3,...$ and $j=\frac{1}{2},\frac{3}{2},\frac{5}{2},...$. Note that energy is proportional to $\chi_{*}$. The groundstate $\left|n=1,\,j=1/2\right\rangle$ and the excited state $\left|n=2,\,j=3/2\right\rangle$ have energy levels
$E_{n=1}^{j=1/2}$ $=\sqrt{1-\alpha^{2}}\,\frac{f\mu}{\sqrt{\lambda}}\,\chi_{*}$ (31)
$E_{n=2}^{j=3/2}$ $=\sqrt{1-\text{\scriptsize$\frac{1}{4}$}\alpha^{2}}\,\frac{f\mu}{\sqrt{\lambda}}\,\chi_{*}$ (32)
A transition from the (initial) excited state $\left|i\right\rangle=\left|n=2,\,j=3/2\right\rangle$ to the (final) groundstate $\left|f\right\rangle=\left|n=1,\,j=1/2\right\rangle$, induced by electric dipole, is allowed as it satisfies the selection rule $\Delta j=\pm 1$. The energy of the photon emitted is $E_{n=2}^{j=3/2}-E_{n=1}^{j=1/2}$, and the frequency of the emitted photon is
$\nu$ $=\frac{E_{n=2}^{j=3/2}-E_{n=1}^{j=1/2}}{2\pi\,\hbar_{*}}$ (33)
$=\frac{f\mu}{2\pi\sqrt{\lambda}}\Bigl{(}\sqrt{1-\text{\scriptsize$\frac{1}{4}$}\alpha^{2}}-\sqrt{1-\alpha^{2}}\Bigr{)}\,\chi_{*}^{3/2}$ (34)
Therefore, the propagation of the photon has the time scale that behaves as
$\tau:=\frac{1}{\nu}\propto\chi_{*}^{-3/2}$ (35)
which is in perfect agreement with Eq. (27).
The time scaling (27) implies that the evolution rate of a clock—regardless of whether it is a mechanical clock, an electronic clock, or an atomic clock—varies in spacetime, as a function of $\chi$, in an anisotropic fashion. In principle, this effect can be measured experimentally: Prepare two identical clocks at a location $A$. Keep one clock at location $A$ and send the second clock to a location $B$. Suppose that the background dilaton field has different values $\chi_{A}$ and $\chi_{B}$ at the two locations $A$ and $B$, respectively. At their respective locations, the clocks would run at different rates per
$\tau_{A}\propto\chi_{A}^{-3/2};\ \ \ \tau_{B}\propto\chi_{B}^{-3/2}$ (36)
When the clock from location $B$ is brought back to location $A$, it will show a different elapsed time compared to the clock that resided at location $A$ during the whole experiment.
We emphasize that this predicted effect is physical, meaning it is, in principle, measurable by comparing the time lapses of two clocks situated at two separate locations with different values of the dilaton field. This time dilation effect differs from the time dilation effect in GR which is associated with the $g_{00}$ component of the spacetime metric, arises from the dependence of the clock rate on the dilaton field $\chi$, viz. Eq. (36). This new phenomenon is distinct and shall be referred to as a “time dilation effect of the Third kind”, to be investigated in future work.  [^8] In addition to the time dilation effect in GR, there is a well-known effect in Special Relativity where two twice-intersecting time-like paths can have different total amounts of proper time in between. Therefore, we refer our predicted time dilation effect as “the Third kind”.
Decay rate of unstable quantum systems
The prediction that we have just made can be validated via a different setup. Reconsider the Hydrogen atom. Induced by electric dipole perturbation in the vacuum, the electron in the excited state $\left|n=2,\,j=3/2\right\rangle$ can spontaneously transition to the groundstate $\left|n=1,\,j=1/2\right\rangle$ (allowed by the selection rule $\Delta j=\pm 1$). According to quantum mechanics [137], the decay rate (i.e. Einstein’s coefficient) is known to be
$A=\frac{4}{3}\frac{\omega_{if}^{3}e^{2}\left\langle\vec{r}_{if}\right\rangle^{2}}{\hbar_{*}c_{*}^{3}}$ (37)
where the effective $\hbar_{*}$ and $c_{*}$ are made explicit. The angular frequency (with $h_{*}:=2\pi\hbar_{*}$) is
$\omega_{if}=\frac{1}{h_{*}}\left(E_{n=2}^{j=3/2}-E_{n=1}^{j=1/2}\right)$ (38)
and the matrix element of the electric dipole between the initial state $\left|i\right\rangle:=\left|n=2,\,j=3/2\right\rangle$ and the final state $\left|f\right\rangle:=\left|n=1,\,j=1/2\right\rangle$ is $\left\langle\vec{r}_{if}\right\rangle:=\left\langle f|\vec{r}|i\right\rangle$. Using $e^{2}=\alpha\,\hbar_{*}c_{*}$, we get
$A=\frac{4\alpha}{3}\frac{m^{3}c_{*}^{4}}{h_{*}^{3}}\left\langle\vec{r}_{if}\right\rangle^{2}$ (39)
Given that $\left\langle\vec{r}_{if}\right\rangle\propto\chi_{*}^{-1}$, $c_{*}\propto\chi_{*}^{1/2}$, $h_{*}\propto\chi_{*}^{-1/2}$, we find
$A\propto\frac{\chi_{*}^{2}}{\chi_{*}^{-3/2}}\,\chi_{*}^{-2}\propto\chi_{*}^{3/2}$ (40)
Thus, the lifetime of the decay process for this unstable quantum system behaves as
$\tau:=\frac{1}{A}\propto\chi_{*}^{-3/2}$ (41)
which perfectly conforms with Eq. (27). Consequently, the clock involved in our experiment proposed earlier in this section can be any generic timekeeping device. It can be a mechanical clock, an electronic clock, an atomic clock, or even a radioactive quantum system.  [^9] The Hafele–Keating experiment [138, 139] was carried out using atomic clocks.
V Comparison of our VSL scheme with the Fujii–Wetterich scheme
We must stress that the dependence of the clock rate on $\chi_{*}$ has been documented in the works of Fujii and Wetterich [20, 14, 15, 16, 17, 18], although it was not a focal point in their analysis; see Footnote 2. However, their findings diverge from ours, particularly in how the clock rate scales with $\chi_{*}$. Specifically, they predicted a relation
$\tau_{\text{FW}}^{-1}\propto\chi_{*}$ (42)
where “FW” refers to Fujii–Wetterich. This contrasts with our prediction $\tau^{-1}\propto\chi_{*}^{3/2}$, given in Eq. (27).
The distinction arises due to a different set of conditions used in Fujii and Wetterich’s treatments [20, 14, 15, 16, 17, 18]. These authors judiciously kept $\hbar$ and $c$ independent of $\chi_{*}$, setting $\hbar_{\text{FW}}=1$ and $c_{\text{FW}}=1$. Instead, they allowed the electron mass $m$ to be a field, while the electron charge $e$ remained a constant. With regard to Eqs. (15)–(16) and Eqs. (17)–(18), their conditions then produce the following identities
$m_{\text{FW}}:=\frac{f\mu}{\sqrt{\lambda}}\,\chi_{*};\ \ \ e:=\sqrt{\alpha}$ (43)
Consequently, Newton’s “constant” varies as $G_{\text{FW}}:=\frac{c_{\text{FW}}^{3}}{\hbar_{\text{FW}}\chi_{*}^{2}}=\chi_{*}^{-2}$. Instead of Eq. (23), the time evolution of the electron wavefunction in the FW scheme is governed by
$i\,\hbar_{FW}\,\frac{\partial}{\partial t}\psi(t)$ $=\hat{H}(t)\,\psi(t)$ (44)
with the Hamiltonian $\hat{H}(t)$ given in Eq. (24). Here, $\hbar_{FW}$ is constant (equal to $1$). Therefore, the time scale for evolution in the FW scheme scales as
$\tau_{\text{FW}}\simeq\frac{\hbar_{\text{FW}}}{\hat{H}}\propto\chi_{*}^{-1}$ (45)
This result can be verified with the concrete examples we used previously. In the FW scheme, the frequency of emitted photon from the excited state $\left|n=2,\,j=3/2\right\rangle$ to the groundstate $\left|n=1,\,j=1/2\right\rangle$ is
$\nu_{\text{FW}}$ $=\frac{E_{n=2}^{j=3/2}-E_{n=1}^{j=1/2}}{2\pi\,\hbar_{\text{FW}}}$ (46)
$=\frac{f\mu}{2\pi\sqrt{\lambda}}\Bigl{(}\sqrt{1-\text{\scriptsize$\frac{1}{4}$}\alpha^{2}}-\sqrt{1-\alpha^{2}}\Bigr{)}\,\chi_{*}$ (47)
instead of the result (34) of our scheme. Likewise, in their scheme, the decay rate of spontaneous emission from the excited state $\left|n=2,\,j=3/2\right\rangle$ to the groundstate $\left|n=1,\,j=1/2\right\rangle$ reads
$A=\frac{4\alpha}{3}\frac{m_{\text{FW}}^{3}\,c_{\text{FW}}^{4}}{h_{\text{FW}}^{3}}\left\langle\vec{r}_{if}\right\rangle^{2}\propto\frac{\chi_{*}^{3}.1^{4}}{1^{3}}\,\chi_{*}^{-2}\propto\chi_{*}$ (48)
instead of Eqs. (39)–(40) of our scheme.
Therefore, despite having an identical matter action, Eq. (14), our scheme and the FW scheme are not physically equivalent. They result in decisively different predictions for the behavior of the clock rate. Future technologies may be able to distinguish the two predictions, and hence the validity of each scheme.
Properties of our mechanism
Our approach offers two distinct benefits:
1. The FW scheme treats inertial mass and gauge charge at disparity: while mass is promoted to fields dependent on $\chi_{*}$, charge remains as a fixed parameter (see Eq. (43)). Given that inertial mass and gauge charge are intrinsic properties of particles, our scheme treats them on equal footing: particle mass and charge are parameters but not fields. (Note: in our scheme, $m$ and $e$ can ‘run’ in the RG flow, but they are independent of the background dilaton $\chi$.)
2. The FW scheme only applies to massive particles, leaving massless particles unaffected. In contrast, the varying $c$ and $\hbar$ in our scheme impact all particles—massive and massless alike. Specifically, the variation in $c$ influences the propagation of (massless) photon in an expanding intergalactic space in cosmology, a feature absent in the FW scheme.
In conclusion of our derivation, the $3/2$–exponent in our time scaling (27) is a novel discovery, leading to new physics with a unique set of previously unexplored phenomena and predictions. Specifically, it induces the variability of the speed of light in spacetime, causing lightwaves traveling through the expanding universe to undergo an additional refraction effect. This effect thence necessitates a reanalysis of the Pantheon Catalog of Type Ia supernovae, a task we carry out in the next section.
VI Phenomenology of VSL
This section focuses on our application of variable speed of light (VSL) for late-time cosmology, bypassing the need for dark energy (DE). The detailed work containing full technicality is presented in Refs. [21].
VI.1 VSL cosmology and late-time cosmography
Our starting point is to generalize the Friedmann–Lemaître–Robertson–Walker (FLRW) metric for the cosmic scale factor (with $\kappa=\{1,0,-1\}$)
$ds^{2}$ $=c^{2}\,dt^{2}-a^{2}(t)\,\biggl{(}\frac{dr^{2}}{1-\kappa\,r^{2}}+r^{2}\,d\Omega^{2}\biggr{)}$ (49)
$d\Omega^{2}$ $=d\theta^{2}+\sin^{2}\theta\,d\phi^{2}$ (50)
by allowing the speed of light $c$ to vary alongside the dilaton field $\chi$ (per $c\propto\chi^{1/2}$ as specified in Eq. (21)). Since the dilaton directly determines the lengthscale for physical processes, $\chi\propto l^{-1}$, as in Eq. (26), and the cosmic scale factor plays the role of a lengthscale in the FLRW metric, it is reasonable to make the following ansatz:
Ansatz #1:
$\chi\propto a^{-1}$ (51)
Furthermore, since the dilaton also determines the timescale for physical processes, $\chi\propto\tau^{-2/3}$, as in Eq. (27), and the cosmic time $t$ plays the role of a timescale in the FLRW metric, it is reasonable to make the following ansatz on the evolution of the cosmic scale factor:
Ansatz #2:
[…]
This ansatz is identical to the evolution of the scale factor in the Einstein–de Sitter (EdS) universe. Obviously, the vanilla EdS model—by itself—cannot account for the late-time cosmic acceleration observed in the Hubble diagram of Type Ia supernova (SNeIa). This failure forms the basis to replace the EdS model by the “concordance” cosmological model in which a $\Lambda$ component of density $\Omega_{\Lambda}\approx 0.7$.
However, in the rest of this paper, we consider the EdS universe in conjunction with a varying speed of light. From $c\propto\chi^{1/2}$ and Ansatz #1, we deduce that, in the intergalactic space, the following relation holds
$c\propto\chi^{1/2}\propto a^{-1/2}$ (53)
The modified FLRW metric suitable for our VSL cosmology is thus expressed as
$ds^{2}=\frac{c_{0}^{2}}{a(t)}\,dt^{2}-a^{2}(t)\,\biggl{(}\frac{dr^{2}}{1-\kappa\,r^{2}}+r^{2}\,d\Omega^{2}\biggr{)}$ (54)
where the cosmic scale factor at our current time $t_{0}$ is set equal to 1 (i.e., $a_{0}=1$), and $c_{0}$ is the speed of light measured in the intergalactic space at $t_{0}$. It is important to note that several authors have previously applied a VSL cosmology to SNeIa standard candles, while circumventing dark energy [28, 34, 31, 35, 33, 32, 36, 29, 30]. The consensus among these studies is that VSL alone could not adequately account for the SNeIa data. However, these conclusions warrant reconsideration, as all previous analyses implicitly assumed the speed of light depends solely on cosmic time $t$. This assumption fails to hold for our VSL cosmology whereby $c$ varies in both space and time, making it location-dependent. This is because the dilaton field $\chi$ can vary in spacetime, with its value in the intergalactic space (which is subject to cosmic expansion) differing from that in gravitationally bound galaxies (which host SNeIa and are resistant to cosmic expansion). Hence, as $c\propto\chi^{1/2}$, $c$ is not only time-dependent but also location-dependent. A thorough treatment of this intricate issue is provided in Ref. [21].
In the presence of VSL, the classic Lemaître redshift formula, $1+z=a^{-1}$, in standard cosmology is no longer applicable. Previous analyses of SNeIa using VSL overlooked this crucial distinction and continued to use the classic formula, resulting in incorrect conclusions. The dependence of $c$ on the dilaton field $\chi$ introduces two fundamental modifications to the right hand side of the Lemaître redshift relation, as follows:
  1. 1.
    The standard $a^{-1}$ term is replaced with $a^{-3/2}$, reflecting the $3/2$–exponent associated with the anisotropic time scaling, Eq. (27).
  2. 2.
    An additional correction accounts for the variation of the dilaton field among galaxies as the function of redshift $z$. This is because galaxies hosting SNeIa may possess different values for the dilaton field due to their evolution in the expanding cosmic background. We model this physically motivated effect by introducing a monotonic function $F(z)$ that smoothly interpolates between two extremes: $F(z\rightarrow 0)=1$ and $F(z\rightarrow\infty)=F_{\infty}$ (with $F_{\infty}$ being an adjustable parameter). Specifically, we find the following formulation to be suitable
    $F(z)=1+(1-F_{\infty})\bigl{(}1-(1+z)^{-2}\bigr{)}^{2}$ (55)
    By virtue of the length scaling (26), $F(z)$ thus signifies the evolution in the typical size of galaxies in response to the expansion of the intergalactic space. This astronomical effect will have implications for the $H_{0}$ tension, which will be shown later in Section VI.8.
Derivations of these modifications are detailed in Ref. [21]. Here, we only quote the results. In our VSL cosmology, the modified Lemaître redshift formula reads
$1+z=a^{-3/2}\,F(z)$ (56)
At very low $z$, with $a(t)=1-H_{0}\,d/c_{0}+\dots$ and $F(z)\approx 1+4(1-F_{\infty})z^{2}+\dots$, the relation (56) yields a modified Hubble law
$z\approx\frac{3}{2}H_{0}\,\frac{d_{L}}{c_{0}}$ (57)
This result contrasts with the Hubble law $z\approx H_{0}\,d/c$in standard cosmology. The multiplicative $3/2$–factor in Eq. (57) will have crucial implications in the estimation of $H_{0}$, as we shall see in Sections VI.3 and VI.5.
Combining Eq. (56) and Ansatz #2 (52), we obtain the modified luminosity distance–redshift relation applicable for our VSL cosmology [21]
$\frac{d_{L}}{c_{0}}=\frac{1+z}{H_{0}\,F(z)}\,\ln\frac{(1+z)^{2/3}}{F(z)}$ (58)
For comparison, this formula fundamentally differs from the canonical ones in the spatially flat $\Lambda$CDM model ($\Omega_{M}+\Omega_{\Lambda}\!=\!1$)
$\frac{d_{L}}{c}=\frac{1+z}{H_{0}}\int_{0}^{z}\frac{dz^{\prime}}{\sqrt{\Omega_{M}(1+z^{\prime})^{3}+\Omega_{\Lambda}}}$ (59)
and in the EdS universe (i.e. $\Omega_{M}=1,\,\Omega_{\Lambda}=0$)
$\frac{d_{L}}{c}=2\,\frac{1+z}{H_{0}}\,\biggl{(}1-\frac{1}{\sqrt{1+z}}\biggr{)}$ (60)
VI.2 Analyzing Pantheon Catalog of SNeIa using VSL
The modified $d_{L}$-vs-$z$ relation (58), supplemented with (55), now stands ready for application to the Pantheon Catalog of SNeIa. Our detailed analysis is provided in Ref. [21]. The Pantheon dataset, as provided in Ref. [140], is shown as open circles in the Hubble diagram in Fig. 1. It comprises $N=1,048$ data points given in terms of redshift and luminosity distance $\{z,\,\mu\}$, together with an error bar $\sigma$ of $\mu$, for $z$ spanning the range (0, 2.26)02.26(0,\,2.26)( 0 , 2.26 ). For a given theoretical model, one would minimize the error
$\mu:=m-M=5\log_{10}(d_{L}/\text{Mpc})+25$ (61)
$\chi^{2}:=\frac{1}{N}\sum_{i=1}^{N}\biggl{(}\frac{\mu_{i}^{\text{model}}-\mu_{i}^{\text{Pantheon}}}{\sigma_{i}^{\text{Pantheon}}}\biggr{)}^{2}$ (62)
Let us first review the canonical interpretation of SNeIa data in standard cosmology. For very low $z$, both formulae (59) and (60) yield $\frac{d_{L}}{c}\approx\frac{z}{H_{0}}$. However, for positive $\Omega_{\Lambda}$, given the same $H_{0}$, the flat $\Lambda$CDM formula (59) produces a higher value for $d_{L}$ compared to the EdS formula (60). The crossover occurs at a redshift $z_{*}$ which approximately satisfies $\Omega_{M}(1+z_{*})^{3}\approx\Omega_{\Lambda}$. We fit Formula (59) to the Pantheon Catalog and obtain $H_{0}=70.2$, $\Omega_{M}=0.285$, $\Omega_{\Lambda}=0.715$ with an error $\chi_{\text{min}}^{2}=0.98824$. Hence, starting at $z\gtrsim z_{*}\approx 0.36$, the spatially flat $\Lambda$CDM model begins to show an excess in $d_{L}$, meaning high-$z$ SNeIa appear dimmer than the EdS model would predict. This behavior, manifest in the high-$z$ portion of the Hubble diagram (Fig. 1), has been interpreted as evidence for late-time cosmic accelerating expansion [141, 142] and for the existence of dark energy.
Our VSL formula (58) involves two adjustable parameters $H_{0}$ and $F_{\infty}$. Its best fit to the Pantheon Catalog yields $H_{0}=47.2$, $F_{\infty}=0.93$ with an error $\chi_{\text{min}}^{2}=0.98556$, which is highly competitive in quality to the $\Lambda$CDM fit. Note that in the upper panel of Fig. 1, the two curves—VSL and $\Lambda$CDM—are indistinguishable in the range $z\lesssim 2.26$ available for the Pantheon data  [^10] Light sources with even higher $z$, such as quasars [143], might help distinguish Formula (58) versus Formula (59)..Both curves manifest an excess in $d_{L}$ in the high-$z$ segment of the Hubble diagram. For completeness, we also show $F(z)$ and $F(a)$ as solid curves in Fig. 3.
Figure 1: Hubble diagram of SNeIa in the Pantheon Catalog. Open circles: 1,048 data points with error bars, listed in Ref. [140]. In both panels, long-dashed line is the $\Lambda$CDM Formula (59) with $H_{0}=70.2$, $\Omega_{M}=0.285$, $\Omega_{\Lambda}=0.715$; dotted line is the EdS Formula (60) with $H_{0}=70.2$. Upper panel: Solid line is our VSL Formula (58) with $H_{0}=47.2$ and $F_{\infty}=0.93$. Lower panel: Solid line is our VSL Formula (58) with $F(z)\equiv 1\ \forall z$ with $H_{0}=44.4$.
Figure 2: Solid curves: $F(z)$ and $F(a)$. Dashed curves: $H_{0}(z)/H_{0}(z\text{=}0)$ and $H_{0}(a)/H_{0}(a=1)$.
Figure 3: Posterior distributions of $H_{0}$ (left panel) and $t_{0}$ (right panel), with 68% CL and 95% CL bands shown.
Of the two modifications to our modified Lemaître redshift formula $1+z=a^{-3/2}\,F(z)$ given in (56), the $a^{-3/2}$ term is expected to be of primary significance, while the $F(z)$ plays a secondary role. To test this intuition, we disable $F(z)$ by setting $F_{\infty}=1$ (making $F(z)\equiv 1\ \forall z\in\mathbb{R}$). A refit of (58) to the Pantheon data with just one parameter (i.e. $H_{0}$) yields $H_{0}=44.4$ with $\chi_{\text{min}}^{2}=1.25366$. Despite a substantial decrease in quality of fit, the $a^{-3/2}$ term alone still produces a reasonable fit, as shown in the lower panel of Fig. 1. Most importantly, this simplified fit continues to exhibit the excess in $d_{L}$ prominent in the high-$z$ segment of the Hubble diagram.
This indicates that our analysis practically—and parsimoniously—requires only one parameter $H_{0}$. The function $F(z)$ plays a role in resolving the $H_{0}$ tension, which will be explained in Section VI.8 below.
VI.3 A reduced value of H0 by a factor of $3/2$
A striking result deduced from our SNeIa analysis thus far is a low value $H_{0}=47.2$, which starkly contrasts with the local measurement $H_{0}=73$ widely accepted in the $\Lambda$CDM framework. Note that this low value is inherent to the modified $d_{L}$-vs-$z$ formula (58) regardless of the function $F(z)$. When $F(z)$ is disabled, the formula still produces $H_{0}=44.4$, a value comparable to that obtained when $F(z)$ is enabled. The posterior distribution of $H_{0}$ is shown in the left panel of Fig. 3 and corresponds to $H_{0}=47.2\pm 0.4$ (at 95% confidence level).
We should note that a method of “redshift remapping” was introduced in [144, 145, 146]. Our modified Lemaître redshift formula could be viewed as effectively a “redshift remapping”. Interestingly, in [146], based on this method, a joint analysis of all primary cosmological probes including the local measurement of the Hubble constant, SNeIa, Baryon Acoustic Oscillations (BAO), Planck observations of the CMB power spectrum and cosmic chronometers yields $H_{0}=48\pm 2$, a range in good agreement with our result $H_{0}=47.2\pm 0.4$.
The drastic deviation in the current-time Hubble value between our $H_{0}=47.2$ from the local measurement $H_{0}=73$ can be understood by using the modified Hubble law (57) in VSL cosmology. For very low $z$, the Hubble law for our VSL cosmology is modified to $z\approx\frac{3}{2}H_{0}\,d/c_{0}$, as opposed to the standard Hubble law $z\approx H_{0}\,d/c$. Thus, for our VSL cosmology, the numerical value (empirically known to be $\sim 71$) for the slope of the $z$-vs-$d$ relation is not $H_{0}$ but $\frac{3}{2}H_{0}$, which promptly leads to $H_{0}\sim 47$. In conclusion, it is the $3/2-$exponent in the $a^{-3/2}$ term in our modified Lemaître redshift formula (56) that is responsible for the reduction in the $H_{0}$ value by a factor of $3/2$ compared to the canonical range.
Surprisingly, this reduced value of $H_{0}$ finds corroboration in a remarkable proposal made by Blanchard et al in 2023 [147], when these authors analyzed the CMB power spectrum. We shall review their work below.
VI.4 Revitalizing Blanchard et al (BDRS)’s analysis of CMB power spectrum, bypassing DE
It is well established that the $\Lambda$CDM, with the primordial fluctuations spectrum $P(k)\propto k^{n}$, successfully accounts for Planck’s thermal power spectrum of the Cosmic Microwave Background (CMB) with parameters $n\approx 0.96$, $\Omega_{\Lambda}\approx 0.7$, $H_{0}\approx 67$. However, in 2003, Blanchard, Douspis, Rowan-Robinson, and Sarkar (BDRS) made a surprising discovery [147]. Instead of adhering to the standard primordial fluctuation spectrum $P(k)\propto k^{n}$, they considered two separate power-law forms for the low $k$ and high $k$ regimes, viz.
$P(k)=\begin{cases}A_{1}\,k^{n_{1}}&\text{for }k\leqslant k_{*}\\ A_{2}\,k^{n_{2}}&\text{for }k\geqslant k_{*}\end{cases}$ (63)
where $k_{*}$ is a break point, and continuity across $k_{*}$ is enforced by imposing $A_{1}\,k_{*}^{n_{1}}=A_{2}\,k_{*}^{n_{2}}$. Most importantly, they deliberately set $\Omega_{\Lambda}=0$, thereby avoiding the dark energy hypothesis. Remarkably, their model achieved an excellent fit to WMAP’s CMB data available at the time, with the optimal parameter values including $H_{0}=46$, $k_{*}=0.0096$ Mpc-1, $n_{1}=1.015$, $n_{2}=0.806$ [147]. Of these parameters, the most striking aspect was the low Hubble value $H_{0}=46$. A subsequent in-depth study by Hunt and Sarkar [148], utilizing a supergravity-based multiple inflation scenario, also successfully accounted for the CMB power spectrum without invoking dark energy, while requiring a similar value of $H_{0}\approx 43.5$. (Interestingly, Shanks in 2004 [149] also ventured that if $H_{0}\lesssim 50$ then a simpler, inflationary model with $\Omega_{baryon}=1$ might be allowed with no need for dark energy or cold dark matter.)
[…]
It is also important to note that the perceived low $H_{0}$ value $\sim 46$ eventually led BDRS to abandon their original findings in a follow-up study related to SDSS’s two-point correlation data of luminous red galaxies [150]. However, their abandonment may have been premature, since in the VSL framework the anomalous $3/2$–exponent in the anisotropic time scaling influences the observation of distant light sources along the time direction. Therefore, a reassessment of BDRS’s work of SDSS data, incorporating this effect, would be warranted. Additionally, future investigations should explore the impacts of VSL on phenomena such as lensing, BAO, and the Etherington distance duality relation.
As we stated earlier in Section VI.3, using a “redshift remapping” method for all primary cosmological probes, the authors in [146] obtained $H_{0}=48\pm 2$, which is compatible with our value $H_{0}=47.2$ and the value $H_{0}=46$ obtained by BDRS.
VI.5 Resolving the age problem, bypassing DE
The spatially flat $\Lambda$CDM model gives the age formula
$t_{0}^{\Lambda\text{CDM}}=\frac{2}{3\sqrt{\Omega_{\Lambda}}H_{0}}\,\text{arcsinh}\sqrt{\frac{\Omega_{\Lambda}}{\Omega_{M}}}$ (64)
With $H_{0}=70.2$, $\Omega_{M}=0.285$, $\Omega_{\Lambda}=0.715$, it yields an age of $13.6$ billion years, an accepted figure in standard cosmology. If dark energy is absent ($\Omega_{\Lambda}=0,\,\Omega_{M}=1$), Eq. (64) simplifies to $t_{0}^{\text{EdS}}=2/(3H_{0})$ which then, with $H_{0}\sim 71$, would result in an age of $9.2$ Gy which would be too short to accommodate the existence of the oldest stars—a paradox commonly referred to as the age problem.
However, our VSL cosmology naturally resolves the age problem without requiring dark energy. Using the definition $H_{0}:=\left.\frac{1}{a}\frac{da}{dt}\right|_{t=t_{0}}$ and the evolution law (52), the age of the our VSL universe is
$t_{0}^{\text{VSL}}=\frac{2}{3H_{0}}$ (65)
[…]
Figure 4: The physical intuition of late-time acceleration based on VSL as explained in Section VI.7. In the left panel, photons from SNe A and SNe C were emitted at times $t_{A}$ and $t_{C}$ with $t_{C}-t_{0}=2\,(t_{A}-t_{0})$; thus, their redshifts satisfy $z_{C}=2\,z_{A}$. However, for the photons emitted from SNe C, the earlier segment of their trajectory had a higher speed of light than the later segment, allowing them to cover a longer distance. This results in $d_{C}>2\,d_{A}$, leading in an excess distance modulus in the Hubble diagram for high-$z$ SNe, as depicted in the right panel.
VI.6 Obtaining late-time acceleration via VSL, bypassing DE
It is an empirical fact that high-$z$ SNeIa’s appear fainter than what would be predicted by the EdS model [141, 142]. This phenomenon is reflected in the distance modulus excess observed in the high-$z$ segment of the Hubble diagram in Fig.  1. Standard cosmology attributes this behavior to late-time cosmic acceleration, necessitating the introduction of a cosmological constant $\Lambda$ into the Friedmann equations, resulting in the $\Lambda$CDM model where $\Omega_{\Lambda}$ quantifies the amount of dark energy.
However, the analysis presented in Sections VI.1 and VI.2 demonstrates the efficacy of our VSL cosmology in accounting for the excess in $d_{L}$ without invoking dark energy. Its success can be quantitatively understood as follows. At high $z$, the luminosity distance formula of the EdS model (60) behaves as
$d_{L}\propto z$ (66)
In contrast, the modified formula for our VSL cosmology (58) behaves as
$d_{L}\simeq z\,\ln z$ (67)
which gains an additional $\ln z$ term compared to (66). Consequently, according to our VSL Formula, $d_{L}$ grows faster and exhibits a steeper upward slope in the high-$z$ section in the Hubble diagram. Furthermore, this explanation can be strengthened through a qualitative account, as we will illustrate below.
VI.7 Physical intuition of late-time accelerationvia VSL, bypassing DE
Consider two supernovae, $A$ and $B$, located at distances $d_{A}$ and $d_{B}$ from the Earth, with $d_{B}=2\,d_{A}$. In standard cosmology, their redshift values $z_{A}$ and $z_{B}$ are related as $z_{A}\approx 2\,z_{A}$ (to first-order approximation). However, this relationship breaks down in our VSL cosmology. Since $c\propto a^{-1/2}$, light traveled faster in the distant past (when the cosmic factor $a\ll 1$) than in the more recent epoch (when $a\lesssim 1$). Therefore, photons emitted from supernova $B$ could cover twice the distance in less than double the time required by photons emitted from supernova $A$. Having spent less time in transit than predicted by standard cosmology, the $B$–photons experience less cosmic expansion than expected, resulting in a lower redshift such that:
$z_{B}<2\,z_{A}\ \ \text{for}\ \ d_{B}=2\,d_{A}$ (68)
Conversely, consider a supernova $C$ with $z_{C}=2\,z_{A}$. For photons emitted from supernova $C$ to experience twice the redshift of the $A$–photons, they must travel a distance greater than twice that of the $A$–photons:
$d_{C}>2\,d_{A}\ \ \text{for}\ \ z_{C}=2\,z_{A}$ (69)
This occurs because the $C$–photons traveled faster at the beginning of their journey toward Earth and must originate from a farther distance (thus appearing fainter than expected) to accumulate sufficient redshift. In either case, the VSL-based Hubble diagram exhibits a steeper upward slope in the high-$z$ section. This intuition is illustrated in Fig. 4.
An alternative explanation for late-time acceleration is a decrease in the speed of light in the intergalactic space as the universe expands. In summary, VSL cosmology can account—qualitatively and quantitatively—for the distance modulus excess observed in high-$z$ SNeIa without the need for dark energy. This approach overcomes the coincidence problem associated with the concordance $\Lambda$CDM model. Furthermore, it bypasses the enigmatic nature of dark energy and aligns with BDRS’s 2003 groundbreaking analysis of the CMB power spectrum which also negates the dark energy hypothesis [147].
VI.8 H0 tension:   A potential resolution with an astronomical origin
In recent years, interest in the $H_{0}$ tension has intensified [1, 2, 151]. One popular approach to address this issue involves treating $H_{0}$ as a “running” value, meaning that its value depends on the redshift of the data used to deduce it; e.g., see Refs. [154, 152, 153].
Our VSL cosmology naturally produces a “running” $H_{0}$ through the function $F(z)$. As mentioned in Section VI.1, $F(z)$ represents the change in the typical size of galaxies as they evolve in an expanding intergalactic space (see also Ref. [21] for clarifications). From the modified formula (58), we can define the “running” value for $H_{0}$ as
$\frac{d_{L}}{c_{0}}=\frac{1+z}{H_{0}(z)}\,\ln(1+z)^{2/3}$ (70)
By comparing this with Eq. (58) itself, we find
$H_{0}(z)=H_{0}\,F(z)\,\biggl{(}1-\frac{3}{2}\frac{\ln F(z)}{\ln(1+z)}\biggr{)}^{-1}$ (71)
If $F(z)\equiv 1\ \forall z>0$, then $H_{0}(z)\equiv H_{0}$, a constant. Otherwise, as $z\rightarrow\infty$, $H_{0}(z)\rightarrow H_{0}F_{\infty}$; namely, at very high $z$, its “running” value is reduced by a factor $F_{\infty}$.
Using (55) and the value $F_{\infty}=0.93$ obtained from the fit in Section VI.2, the computed “running” $H_{0}(z)$ is depicted as dashed curves in Fig. 3. At the time of recombination, the “running” value for $H_{0}$ would be $H_{0}\,F_{\infty}=47.2\times 0.93=43.9$, representing a $7\%$ reduction from its value estimated using low-redshift SNeIa. The function $F(z)$ hence shed important light on the $H_{0}$ tension. Furthermore, in [148] Hunt and Sarkar obtained $H_{0}\approx 44$ from the CMB power spectrum using a supergravity-based multiple inflation scenario. Our result of $H_{0}(z\gg 1)\approx 43.9$ is thus in agreement with that of the Hunt–Sarkar team.
Importantly, the function $F(z)$ is a well-defined and physically motivated concept: it models the evolution of galaxies in an expanding intergalactic space. Consequently, the “running” $H_{0}$ in our approach is rooted in an astronomical origin, rather than a cosmic origin.
VI.9 An infinite cosmological horizon using VSL
In the modified FLRW metric (49), the cosmological horizon at the recombination time $t_{rec}$ is given by
[…]
Since $c=c_{0}\,a^{-1/2}$ and $a=\left(t/t_{0}\right)^{2/3}$, we find
$l_{H}\left(t_{rec}\right)$ $=c_{0}\,t_{0}\,\left(\frac{t_{rec}}{t_{0}}\right)^{2/3}\int_{0}^{t_{rec}}\frac{dt^{\prime}}{t^{\prime}}=\infty$ (73)
The cosmological horizon is thus (logarithmically) divergent, suggesting that the entire universe was causally connected at the recombination event. This result dovetails with the VSL proposals previously made by Moffat [26] and Albrecht and Magueijo [27] which sought to explain the near isotropy of the CMB temperatures—an empirical fact commonly known as the horizon problem.
VII Discussion and Summary
On scale symmetry—Symmetries plays a fundamental role in constructing modern physical theories, with notable examples including general covariance in GR and gauge invariance in the Glashow–Weinberg–Salam (GWS) model of particle physics. In his seminal 1995 report [155], Bardeen proposed employing scale symmetry as a viable alternative to supersymmetry to cure the naturalness problem of the Higgs mass. It is well established that the Higgs particles, as quanta of scalar fields, suffer a quadratic divergence in their loop amplitudes, which destabilizes their ultraviolet-limit behavior and requires fine tuning. Bardeen suggested that classical scale invariance could protect the Higgs mass from this divergence. Theories that impose scale invariance in the matter and gravity sectors have been actively pursued in both particle physics and gravitational physics [332, 298, 299, 300, 301, 274, 275, 276, 277, 228, 190, 189, 259, 247, 254, 327, 357, 353, 349, 258, 330, 162, 340, 333, 347, 348, 356, 311, 233, 260, 170, 307, 215, 238, 239, 249, 313, 303, 261, 278, 248, 335, 358, 337, 183, 309, 184, 186, 187, 185, 262, 263, 241, 328, 329, 235, 237, 234, 236, 346, 192, 191, 193, 194, 304, 217, 218, 219, 221, 222, 220, 314, 321, 176, 160, 202, 204, 207, 206, 201, 200, 203, 205, 280, 334, 316, 319, 326, 264, 265, 281, 282, 284, 283, 240, 250, 294, 230, 231, 285, 214, 352, 188, 227, 255, 266, 232, 286, 226, 224, 225, 315, 287, 310, 308, 267, 268, 177, 178, 180, 179, 181, 182, 288, 289, 290, 164, 169, 168, 167, 165, 166, 242, 243, 244, 245, 246, 223, 351, 251, 252, 253, 355, 197, 229, 302, 256, 171, 322, 323, 325, 324, 318, 345, 291, 216, 295, 163, 359, 211, 212, 213, 341, 279, 342, 343, 269, 336, 320, 270, 271, 312, 195, 196, 257, 317, 161, 172, 344, 199, 198, 272, 157, 156, 158, 159, 338, 339, 297, 354, 305, 306, 209, 210, 350, 292, 273, 331, 208, 173, 174, 175, 293, 296, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129]. Scale invariance typically necessitates (non-minimal) coupling of the Higgs field with gravity via a scalar field $\chi$—commonly referred to as a “dilaton”—which arises naturally in various theoretical contexts, such as Kaluza–Klein, string theory, and braneworld scenarios [12].
In Section II, we consider a scale-invariant action of gravity and matter, as described by Eqs. (7), (8), and (12) with $V(\chi)\propto\chi^{4}$. This action is generally covariant, locally Lorentz invariant, and $U(1)$ gauge invariant (for the matter sector). Notably, the scale invariance of the action allows it to evade the observational bounds on the fifth force [123, 115]. The Higgs field couples with the dilaton field $\chi$ in the form $\chi^{2}\,\Phi^{2}$. In addition, the matter Lagrangian $\mathcal{L}_{\text{mat}}$, given in Eq. (8) as a prototype, is invariant with respect to a discrete $\mathbb{Z}_{2}$ symmetry of the Higgs field, viz. $\Phi\leftrightarrow-\Phi$, although it can be readily generalized to encompass the $SU(3)\times SU(2)\times U(1)$ gauge group of the GWS model.
Our mechanism to generate variable $\hbar$ and $c$—Because of scale invariance, all parameters in the action are dimensionless. This crucial fact means that the Planck constant $\hbar$ and the speed of light $c$—as dimensionful quantities—must be absent from the outset of this adimensional action.
In Section III, we developed—for the first time—a recipe to construct $\hbar$ and $c$ from the Higgs-dilaton coupling in the matter sector $\mathcal{L}_{\text{mat}}$ of Eq. (8). The metric $g_{\mu\nu}$ and the dilaton $\chi$ are treated as slowly varying background fields. Within a local set surrounding a given point $x^{*}$, the discrete $\mathbb{Z}_{2}$ symmetry of the Higgs field is spontaneously broken due to the quartic “potential” $\frac{1}{2}\mu^{2}\,\chi^{2}\,\Phi^{2}-\frac{1}{4}\lambda\,\Phi^{4}$ (for $\lambda>0$). The Higgs field then acquires a non-zero VEV $\left\langle\Phi(x^{*})\right\rangle$ which is proportional to $\chi(x^{*})$; see Eq. (13). The non-zero Higgs VEV gives rise to mass for both the fermion and the Higgs scalar boson (while the gauge vector boson remains massless since the $U(1)$ gauge remains unbroken). Importantly, the Higgs VEV allows us to construct a Planck constant $\hbar_{*}$ and a speed of light $c_{*}$ to be effective within the local set around $x^{*}$.
It is worth noting that in Ref. [135], we presented a “short-cut” approach. Instead of invoking the Higgs-dilaton coupling $\chi^{2}\,\Phi^{2}$ and the Yukawa coupling between the fermion and the Higgs field $\bar{\psi}\psi\,\Phi$ as we do in Eq. (8) in this current paper, therein we intentionally suppressed the Higgs field and let the fermion directly couple with the dilaton via the term $\chi\,\bar{\psi}\psi$. In doing so, we also obtained a construction for $\hbar_{*}$ and $c_{*}$ similar to that presented in this paper. However, the current paper aims to explicitly incorporate the Higgs field and utilize the mechanism of spontaneous symmetry breaking. so that a connection to the GWS model for the matter sector can be established.
In either approach, the end result is that each open set vicinity of a given point $x^{*}$ on the spacetime manifold is equipped with a replica of the Quantum Electrodynamics (QED) action for matter fields (i.e. fermion, $U(1)$ gauge vector boson, and Higgs scalar boson), operating with its own effective values of $\hbar_{*}$ and $c_{*}$. Lorentz symmetry is valid locally within each open set. These QED replicas are isomorphic (i.e. structurally identical), differing only in the values of $\hbar_{*}$ and $c_{*}$, which are determined by the prevailing value of the background dilaton field $\chi(x^{*})$ per $\hbar_{*}=\chi^{-1/2}(x^{*})$ and $c_{*}=\chi^{1/2}(x^{*})$. With the gravitational Lagrangian $\mathcal{L}_{\text{grav}}$ in Eq. (12) governing the dynamics of $\chi$, the values of $\hbar_{*}$ and $c_{*}$ thus vary from one open set to another across spacetime.
Crucially, within each open set, the length and time scales of physical processes are governed by the dilaton field as $l\propto\chi^{-1}(x^{*})$ and $\tau\propto\chi^{-3/2}(x^{*})$. This leads to an anisotropic relationship between the rate of clocks and the length of measuring rods, per $\tau^{-1}\propto l^{-3/2}$. This novel relationship allows us to predict a new “time dilation effect of the Third kind”, to be reported in future work.
The role of VSL in late-time cosmologyThe anisotropic relationship $\tau^{-1}\propto l^{-3/2}$ has immediate cosmological consequences. Section VI presents key findings regarding the impacts of variable speed of light (VSL) on the Hubble diagram of SNeIa, with full technical details of our analysis provided in a companion paper, Ref. [21].
In the presence of VSL, the standard cosmography is no longer applicable. Section VI.1 produces a new cosmography that accommodates VSL:
(i) Under the assumption that the dilaton is related to the cosmic scale factor as $\chi\propto a^{-1}$, the FLRW metric is modified to
$ds^{2}=\frac{c_{0}^{2}}{a(t)}\,dt^{2}-a^{2}(t)\,\left[dr^{2}+r^{2}d\Omega^{2}\right]$
which leads to a modified Lemaître redshift formula (where the function $F(z)$ accounts for variations in the dilaton value among individual galaxies)
$1+z=a^{-3/2}\,F^{3/2}(z)$
and a modified Hubble law
$z=\frac{3}{2}H_{0}\,\frac{d_{L}}{c_{0}}$
[…]
$\frac{d_{L}}{c_{0}}=\frac{1+z}{H_{0}\,F(z)}\,\ln\frac{(1+z)^{2/3}}{F(z)}$
Based on this formula, in Section VI.2, we (re)analyzed the Pantheon Catalog of SNeIa. The key findings are as follows:
* We achieved a fit to the Pantheon data that surpasses the quality of the $\Lambda$CDM model. The function $F(z)$ is parametrized as $F(z)=1-(1-F_{\infty}).\left(1-(1+z)^{-2}\right)^{2}$ with $F_{\infty}=0.931\pm 0.11$ (95% CL).
* We obtained a new Hubble value $H_{0}=47.2\pm 0.4$ (95% CL), which is approximately $2/3$ of the $H_{0}\approx 70$ value derived from SNeIa when relying on the $\Lambda$CDM model (with $\Omega_{\Lambda}\approx 0.7$). The reduction in the $H_{0}$ value arises from the $3/2$–multiplicative factor in the modified Hubble law, Eq. (57). See Section VI.3.
* The reduced value of $H_{0}=47.2$ yields a value of $t_{0}=13.9\pm 0.1$ Gy (95% CL) for the age of an EdS universe, effectively resolving the age paradox without requiring dark energy. See Section VI.5.
* Additionally, the function $F(z)$ offers a potential resolution of the $H_{0}$ tension. See Section VI.8.
Importantly, Sections VI.6 and VI.7 provide a new perspective: the Hubble diagram of SNeIa does not necessarily indicate an accelerating universe. Instead, it supports an alternative interpretation in favor of an EdS universe characterized by a declining speed of light.
On the Blanchard–Douspis–Rowan-Robinson–Sarkar (BDRS) analysis of the CMB power spectrum, avoiding dark energy—Section VI.4 revisits an important—but long-overlooked—proposal from 2003 [147], where BDRS shed new light on the CMB power spectrum. By adopting a double-power form for the primordial fluctuation spectrum as given Eq.  (63), they achieved an excellent fit to WMAP’s CMB power spectrum within an EdS model, namely, without invoking dark energy. A surprising outcome of their analysis was a value of $H_{0}\approx 46$, which, while at odds with the canonical range of $H_{0}\sim 67-73$, aligns remarkably well with the $H_{0}=47.2$ value derived from our analysis of SNeIa. The nearly perfect agreement in the values of $H_{0}$, obtained from two datasets with different natures and covering separate epochs, indicates a consistent cosmological framework based on the EdS model with a variable speed of light, while eliminating the need for dark energy.
VIII Conclusion
From the Higgs-dilaton coupling, we have derived variations in Planck’s quantum of action and the speed of light in curved spacetime. Our work hence—for the first time—fulfills Einstein’s original vision, dating back to 1911-1912 [23, 24, 25], by integrating the local validity of Lorentz symmetry, the variability of the speed of light, and the (Riemannian) geometric nature of gravity into a unified framework.
Importantly, the variable speed of light provides an alternative explanation for the Hubble diagram of SNeIa: Rather than being attributed to a late-time cosmic acceleration, the excess in the distance modulus observed in high-redshift SNeIa is a result of a declining speed of light in an expanding Einstein–de Sitter universe.
Strikingly, the reduced value of $H_{0}=47.2$ derived herein from the SNeIa data aligns remarkably well with the $H_{0}\approx 46$ value that BDRS deduced from the CMB data, all while bypassing the need for dark energy [147]. Taken together, our work and that of BDRS diminish the necessity for the Dark Energy hypothesis, along with its associated fine-tuning and coincidence problems.
In addition, potential implications of scale invariance and our mechanism in realms beyond late-time cosmology are briefly outlined in Appendix A.
Lastly, as $\hbar$ and $c$ become variable in scale-invariant gravity, they can no longer serve as “fundamental units” within Planck’s $cGh$ system [7, 8, 9]. Appendix B introduces a new unit system suitable for variable $\hbar$ and $c$.
Acknowledgements.
I thank Clifford Burgess, Tiberiu Harko, Robert Mann, Anne-Christine Davis, Eoin Ó Colgáin, Subir Sarkar, Leandros Perivolaropoulos, Demosthenes Kazanas, John Moffat, Alberto Salvio, Ilya L. Shapiro, Gregory Volovik and Bruce Bassett for their constructive and insightful feedback. I wish to thank the anonymous Referee for his or her comments, which significantly helped improve the paper.
————$\infty$—————
Appendix A   AN OUTLOOK BEYOND LATE-TIME COSMOLOGY
Here, we speculate on the potential ramifications of scale invariance, as well as variable $c$ and $\hbar$, in realms beyond late-time cosmology.
* Early-time cosmology: Section VI.9 “An infinite cosmological horizon using VSL” explicitly shows that an EdS universe with declining speed of light exhibits an infinite cosmological horizon; see Eq. (73). This result could help resolve the horizon paradox, which was one of the key motivations for the VSL proposals made by Moffat and the Albrecht–Magueijo team in the 1990s [26, 27].
* On the flattening galactic rotation curves: For $\mathcal{L}_{\text{grav}}$ in Eq. (12), the case of $\omega=0$ and $V(\chi)\propto\chi^{4}$ corresponds to quadratic gravity. It has been shown [128, 124] that when expanding around a de Sitter background metric (rather than the flat background conventionally used), the massless rank-2 tensor graviton mode produces a long-range gravitational potential with the correct Newtonian tail, viz. $\sim 1/r$. Furthermore, it has been shown [124] that the massless rank-0 scalar graviton mode yields a long-range potential which grows linearly with distance, viz. $\sim r$. A superposition of these modes could give rise to the Mannheim–Kazanas potential (previously derived for conformal gravity [360, 361]), $-GM/r+\gamma\,r$, where $GM$ and $\gamma$ are two adjustable parameters. In [363, 362], this potential has been employed to successfully account for the observed flattening of rotation curves across a wide range of galaxies, without invoking dark matter.
* Variable $\hbar$: A Planck constant that varies in spacetime could, in principle, alter—both qualitatively and quantitatively—the radiative behavior of black holes in scale-invariant gravity.
* On the Hierarchy Problem and the Cosmological Constant Problem: In quantum field theories of matter in flat spacetime, the loop corrections to the Higgs mass and vacuum energy manifest as quadratic and quartic divergences, respectively [366, 367, 365, 364, 368]. However, it is plausible to suspect that these divergences only appear problematic when considered in the context of flat spacetime—namely, in isolation from the otherwise curved background spacetime. That is to say, these loop corrections currently account for contributions solely from the matter sector, while neglecting those from the gravitational sector. One intuitive way to understand this is that in scale-invariant gravity, as $\hbar$ and $c$ become dependent on the dilaton per $\hbar\propto\chi^{-1/2}$ and $c\propto\chi^{1/2}$, the dilaton—as a component of the gravitational sector—could, in turn, modulate these loop corrections and keep them in check. If so, scale invariance, as inspired by Bardeen’s seminal idea [155], might open a pathway toward protecting the Higgs mass and vacuum energy from divergence and fine tuning.
These intriguing possibilities, albeit speculative at this stage, represent open avenues for future investigation.
Appendix B   ON PLANCK’S AND HARTREE’S UNIT SYSTEMS
In our mechanism, the electron mass $m$ and charge $e$ are dimensionless quantities derived from the parameters $f$, $\mu$, $\lambda$, and $\alpha$ via Eq. (19). The dilaton field $\chi$ is the single entity that carries a scale, which then serves as the only unit available. Consequently, all physical quantities can be expressed exclusively in terms of the magnitude of $\chi$. For instance, $\hbar$ and $c$ have the units of $\chi^{-1/2}$ and $\chi^{1/2}$, respectively, whereas $G$ is dimensionless. The rest mass energy $E=mc^{2}$ has the unit of $\chi$, owing to $m$ being dimensionless and $c\propto\chi^{1/2}$. Similarly, the Coulomb potential of a point charge $e$, given by $e^{2}/r$, also carries the unit of $\chi$, by virtue of $e$ being dimensionless and $r\propto\chi^{-1}$. In general, a theory with dilatation symmetry would require only one ‘unit’ (associated with the dilaton field $\chi$), rendering the canonical three-unit system (mass $M$, length $L$, time $T$) unnecessary. It is worth noting that this conclusion agrees with Ref. [369], which deduced from a spacetime perspective that the number of fundamental constants equals one in relativistic spacetimes.
Nevertheless, to distinguish between mass and charge, which represent two distinct aspects of the electron (viz., inertia and $U(1)$ gauge coupling strength), we introduce two ‘labels’ $m_{0}$ and $e_{0}$ into their definitions, Eq. (19)
$m:=\frac{f\mu}{\sqrt{\lambda}}\,m_{0};\ \ \ e:=\sqrt{\alpha}\,e_{0}$ (74)
This redefinition leads to
$\hbar_{*}:=e_{0}\sqrt{m_{0}}\,\chi_{*}^{-1/2};\ \ c_{*}:=\frac{e_{0}}{\sqrt{m_{0}}}\,\chi_{*}^{1/2};\ \ G:=\frac{e_{0}^{2}}{m_{0}^{2}}$ (75)
Our new unit system based on the trio $\{e_{0},\,m_{0},\,\chi_{*}\}$ deviates from the Planck $cGh$ unit system, which relies on the trio of constants $\{c,\,G,\,\hbar\}$. Note that our new unit system is akin to the unit system introduced by Hartree in 1928 [370], based on the electron charge $e$, mass $m$, and the Bohr radius $a_{B}$ (the typical size of the hydrogen atom in its ground state). In the Hartree system, the Planck constant and speed of light are defined via $\{e,\,m,\,a_{B}\}$ as
$\hbar$ $=e\sqrt{m\,a_{B}};\ \ \ c=\frac{e}{\alpha\sqrt{ma_{B}}}$ (76)
in close resemblance to the first two identities in Eq. (75).